### v1

# Additivity of Crossing Number via Restricted Reidemeister Moves

Abstract
We define a set of restricted Reidemeister moves and show that if $K$ is obtained from $K_{0}\mathbin{\#}K_{1}$ using those moves, then the crossing number of $K$ is at least $c(K_{0})+c(K_{1})$. We also explore topological interpretations of this result.
1 Introduction
It is a long-standing open question whether the minimal crossing number of a composite knot is equal to the sum of the minimal crossing numbers of its factors. The best result for general knots so far has been due to [Lackenby2009] showing that for any finite number of knots $K_{1},\dots,K_{n}$ the inequality $c(K_{1})+\cdots+c(K_{n})\leqslant 152c(K_{1}\mathbin{\#}\cdots\mathbin{\#}K_{n})$ holds. It has also been established that for a certain class of links, $c(L_{1}\mathbin{\#}L_{2})\geqslant c(L_{1})+3$ [Diao2004]. Equality $c(K_{1}\mathbin{\#}K_{2})=c(K_{1})+c(K_{2})$ has been previously established when $K_{1}$ and $K_{2}$ are both alternating knots [Kauffman1987] or both torus knots [Diao2004].
Given a knot diagram $K$, denote by $\chi(K)$ the number of crossings in that diagram. Denote by $\mathcal{R}$ the set of classical Reidemeister moves and Reidemeister equivalence between diagrams as $K\sim_{\mathcal{R}}K^{\prime}$. Then $c(K)$ is defined as the minimim of $\{\chi(K^{\prime})\,:\,K^{\prime}\sim_{\mathcal{R}}K\}$. Now the question of additivity can be reformulated as follows. Is it possible that $K^{\prime}\sim_{\mathcal{R}}K_{0}\mathbin{\#}K_{1}$ and $\chi(K^{\prime})<c(K_{0})+c(K_{1})$? We will define a set of restricted Reidemeister moves $\mathcal{R}^{\prime}$ (in fact several versions) and show that if $K^{\prime}\sim_{\mathcal{R}^{\prime}}K$, then $c(K^{\prime})\geqslant c(K_{0})+c(K_{1})$.
A key part of the proof is bounding the number of certain type of crossings (which we call hybrid crossings) using a graph-theoretic lemma in Section 6.1.
Acknowledgements
The author was generously supported by a European Research Council Advanced Grant (ERC AdG, ILLUSIVE: Foundations of Perception Engineering, 101020977). The author is grateful to Steven LaValle for encouragement and support, Kalle Timperi for early discussions, and Tapani Hyttinen for comments on the proof. Special thanks are due to Marc Lackenby for insightful suggestions and comments on the manuscript.
2 Preliminaries
In this section we will define basic notation used in the paper and the fundamental concepts of genericity and general position, knot and link diagrams, centered and partial diagrams, restricted Reidemeister moves and equivalence, and crossing number.
2.1 Basics
[…]
By $\operatorname{dom}(f)$ denote the domain of a function $f$ and by $|f|$ its range. Often, however, we identify a function $f$ with its range $|f|$ and might denote, for example $f\cap g\subset A$ instead of $|f|\cap|g|\subset A$, or $f\!\restriction\!C\subset g$ instead of $|f\!\restriction\!C|\subset|g|$. It should always be clear from the context when this is the case. If $f$ and $g$ are maps which agree on $\operatorname{dom}(f)\cap\operatorname{dom}(g)$, denote by $f\cup g$ the map with domain $\operatorname{dom}(f)\cup\operatorname{dom}(g)$ such that $(f\cup g)(x)=f(x)$ for $x\in\operatorname{dom}(f)$ and $(f\cup g)(x)=g(x)$ for $x\in\operatorname{dom}(g)$. Sometimes, by abuse of notation, we write the concatenation of curves using $\cup$ instead of , when the intended meaning is clear from the context. The notation $C=A\sqcup B$ means that $C=A\cup B$ and $A\cap B=\varnothing$.
The composition of functions is denoted $f\circ g$ but sometimes $\circ$” is ignored and it is denoted fgfg. If AA is a set, denote by $fA=f[A]=f(A)=|f\!\restriction\!A|=\{f(x)\,:\,x\in A\}$ the image of AA (usually $f[A]$ is preferred). By $\operatorname{id}_{A}$ we denote the identity map $A\to A$. The identity map fixes AA pointwise. Sometimes we say that a function $f$ fixes a set AA pointwise to mean that $f\!\restriction\!A=\operatorname{id}_{A}$.
By $\overset{\circ}{U}{}$, $\partial U$, and $\bar{U}$ denote the interior, closure, and the boundary of a set $U\subset\mathbb{R}^{n}$ respectively. By $\operatorname{Card}(A)$ denote the cardinality of the set AA. If $\gamma_{1}$ and $\gamma_{2}$ are mappings from II into some space $X$ such that $\gamma_{1}(1)=\gamma_{2}(0)$, we denote by $\gamma_{1}{}^{\frown}\gamma_{2}$ the map $\gamma\colon I\to\mathbb{R}^{2}$ given by $\gamma(t)=\gamma_{1}(2t)$ if $t<1/2$, and $\gamma(t)=\gamma_{2}(2t-1)$ otherwise. By $B(x,r)$ and $\overset{\circ}{B}{}(x,r)$ denote the closed and open balls respectively centered at $x$ with radius $r$.
2.2 Genericity
Definition 2.1.
A polyhedron in $\mathbb{R}^{n}$ is a set which is the realization of a finite simplicial complex in $\mathbb{R}^{n}$ compatible with the linear structure of $\mathbb{R}^{n}$. A map $f\colon X\to Y$ between polyhedra $X$ and $Y$ is an immersion if for all $x\in X$ there is a polyhedral neighborhood $U$ of $x$ (i.e. $U$ is a polyhedron and $x\in\overset{\circ}{U}{}$) such that $f\!\restriction\!U$ is an embedding. A point $x\in|f|$ is a multiplicity point if $\operatorname{Card}(f^{-1}(x))>1$. It is a double point if $\operatorname{Card}(f^{-1}(x))=2$. Suppose $T\colon D\to\mathbb{R}^{2}$ is an immersion where $D$ is a 11-manifold with boundary. Suppose that $x\in\mathbb{R}^{2}$ is a double point, Then $x$ is transversal if there is a compact polyhedral neighborhood $U$ of $x$, a homeomorphism $h\colon U\to[-1,1]^{2}$, and disjoint closed connected sets $C_{1},C_{2}\subset D$ each of which contains one of the points of $T^{-1}(x)$ in their interior such that $|h\circ(T\!\restriction\!C_{1})|=[-1,1]\times\{0\}$ and $|h\circ(T\!\restriction\!C_{2})|=\{0\}\times[-1,1]$.
Definition $2.$
Let $D$ a 11-manifold with boundary. We say that an immersion $g$ of $D$ into $\mathbb{R}^{2}$ is generic if there are at most finitely many multiplicity points, all of them are transversal double points and none of them is at $g[\partial D]$. These double points are called crossings and the set of crossings is denoted
$\operatorname{Cross}(g)=\{x\in\mathbb{R}^{2}\,:\,x\text{ is a double point of }g\}.$
2.3 Centered Knot Diagrams
(a) Another knot t-diagram $K$.
(b) The upper partial knot diagram $K^{+}$ of the knot in (a).
(c) The lower partial knot diagram $K^{-}$ of the knot in (a)
Figure 1: (a): A centered knot diagram $K$ and the corresponding upper (b) and lower (c) partial knot diagrams $K^{+}$ and $K^{-}$. Conversely, the diagram
A knot diagram (k.d.) is a pair $(K,\ell)$ such that $K\colon I\to\square$ is a generic immersion with $K(0)=o^{-},K(1)=o^{+}$, $K(t)\in\overset{\circ}{\square}{}$ for all $0<t<1$, and such that the one-sided derivatives of $K$ at $t=0$ and $t=1$ are vertical, and $\ell\colon\operatorname{Cross}(K)\to I^{2}$ is such that $\ell(x)=(a,b)$ where $\{a,b\}=K^{-1}(x)$ for all $x$ (it is an ordering of the inverse image of $x$ under $K$). The diagram is centered, if also $K(\tfrac{1}{2})=o$, $o\notin\operatorname{Cross}(K)$, and the derivative of $K$ at $t=\tfrac{1}{2}$ exists and is vertical. Note that for a piecewise linear map the derivatives exist for all but finitely many points and the one-sided derivatives at the end-points always exist. Denote
$K^{\pm}=K\!\restriction\!I^{\pm}.$ (2.1)
Conventionally in illustrations, $K^{+}$ will be drawn with red color and $K^{-}$ with blue. See Figure 1. More generally, let us define partial knot diagrams:
2.4 Partial Knot Diagrams
An upper (lower) partial knot diagram is a pair $(P,\ell)$ (resp. $(Q,\ell)$) where $P$ (resp. $Q$) is a generic immersion $P\colon I^{+}\to\square$ (resp. $Q\colon I^{-}\to\square$) such that $P(1)=o^{+}$ (resp. $Q(0)=o^{-}$) and the one-sided derivative at $o^{+}$ (resp. $o^{-}$) is vertical. The function $\ell\colon\operatorname{Cross}(P)\to(I^{+})^{2}$ (resp. $\ell\colon\operatorname{Cross}(Q)\to(I^{-})^{2}$) is a crossing labeling function just like for knot diagrams. The partial diagram is centered, if also $P(\tfrac{1}{2})=o$ (resp. $Q(\tfrac{1}{2})=o$) and the one-sided derivative at $\tfrac{1}{2}$ is vertical. Clearly, if $K$ is a centered knot diagram, then $K^{\pm}$ are upper and lower centered partial knot diagrams respectively.
Remark 2.3.
The concept of partial knot diagrams is reminiscient of knotoids introduced in [Turaev2012Knotoids]. The author wishes to thank Prof. Marc Lackenby for pointing this out.
2.5 Restricted Reidemeister Moves
Figure 2: Reidemeister moves and sliding moves
The set of (classical) Reidemeister moves is denoted $\mathcal{R}$. It contains the moves $R_{0},R_{1},R_{2},R_{3}$. The moves $R_{i}$ for $i\in\{1,2,3\}$ are depicted in the top half of Figure 2. The move $R_{0}$ is an application of a self-homeomorphism of $\square$ which is the identity on a neighbourhood of $\partial\square\cup\{o\}$. The equivalence relation “the knot diagram $K$ is obtained from the knot diagram $K^{\prime}$ using Reidemeister moves” is denoted $K\sim_{\mathcal{R}}K^{\prime}$. The corresponding restricted Reidemeister moves $R^{*}_{i}$ are such that $R^{*}_{0}=R_{0}$, and for $i\in\{1,2,3\}$ the move $R^{*}_{i}$ is like $R_{i}$ but has to be performed in a neighborhood which does not contain the critical point $o$. Let $\mathcal{R}^{*}=\{R^{*}_{0},R^{*}_{1},R^{*}_{2},R^{*}_{3}\}$ be the set of non-sliding restricted Reidemeister moves. The equivalence relation corresponding to $\mathcal{R}^{*}$ is denoted $\sim^{*}_{\mathcal{R}}$.
The sliding moves $^{*}_{++}, RR^{*}_{--}$, $R^{*}_{+-}$, and $R^{*}_{-+}$ are moves in which an arc slides under or over the critical point (Figure 2 below). For $\pm\in\{-+\}$, the move $R^{*}_{\pm-}$ allows a subarc of $K^{\pm}$ to slide under the critical point, and the move $R^{*}_{\pm+}$ allows a subarc of $K^{\pm}$ to slide over the critical point.
We define four (extended) sets of restricted Reidemeister moves:
$\mathcal{R}^{*}_{opp1}$ $=\mathcal{R}^{*}\cup\{R^{*}_{++},R^{*}_{--}\}$ $\mathcal{R}^{*}_{same1}=\mathcal{R}^{*}\cup\{R^{*}_{+-},R^{*}_{--}\}$
$\mathcal{R}^{*}_{opp2}$ $=\mathcal{R}^{*}\cup\{R^{*}_{+-},R^{*}_{-+}\}$ $\mathcal{R}^{*}_{same2}=\mathcal{R}^{*}\cup\{R^{*}_{++},R^{*}_{-+}\}$
In the sets $\mathcal{R}^{*}_{opp1}$ and $\mathcal{R}^{*}_{opp2}$, the red and blue arcs (corresponding to $K^{+}$ and $K^{-}$ respectively) are allowed to slide on the opposite sides of the critical point. In $\mathcal{R}^{*}_{opp1}$ the red arc slides over and the blue arc under, and in $\mathcal{R}^{*}_{opp2}$ vice versa. In $\mathcal{R}^{*}_{same1}$ and $\mathcal{R}^{*}_{same2}$ they slide on the same side of the critical point. In $\mathcal{R}^{*}_{same1}$ both slide under and in $\mathcal{R}^{*}_{same2}$ both slide over. The corresponding equivalence relations are respectively denoted
$\sim^{o1}_{\mathcal{R}},\quad\sim^{o2}_{\mathcal{R}},\quad\sim^{s1}_{\mathcal{R}},\quad\text{and}\quad\sim^{s2}_{\mathcal{R}}.$
Remark 2.4.
For $X\in\{o1,o2,s1,s2\}$ and any knot diagrams $K$ and $K^{\prime}$ we have
$K\sim^{*}_{\mathcal{R}}K^{\prime}\Rightarrow K\sim^{X}_{\mathcal{R}}K^{\prime}\Rightarrow K\sim_{\mathcal{R}}K^{\prime}.$
A mirror image of a knot diagram $K=(K,\ell)$ is $\mu(K)=(K,\ell_{\mu})$ where $\ell_{\mu}$ flips every label given by $\ell$, i.e. if $\ell(x)=(a,b)$, then $\ell_{\mu}=(b,a)$. It is now clear that for all knot diagrams $K$ and $K^{\prime}$,
$K\sim^{o1}_{\mathcal{R}}K^{\prime}\iff\mu(K)\sim^{o2}_{\mathcal{R}}\mu(K^{\prime})\qquad\text{ and }\qquad K\sim^{s1}_{\mathcal{R}}K^{\prime}\iff\mu(K)\sim^{s2}_{\mathcal{R}}\mu(K^{\prime}).$
2.6 Composite Knot Diagrams
Let $H=I\times\left\{\tfrac{1}{2}\right\}$ be the horizontal line passing through the middle of $\square$. Suppose that $K$ is a centered knot diagram. If $K\cap H=\{o\}$, we say that $K$ is a composite knot diagram. For $\pm\in\{-,+\}$, let $h^{\pm}\colon I^{\pm}\to I$ be the (unique) linear orientation preserving homeomorphism. Let $\widehat{h}^{\pm}\colon\square^{\pm}\to\square$ be given by $\widehat{h}^{\pm}(x,y)=(x,h^{\pm}(y))$. Further, let $h^{\pm}_{*}\colon I^{2}\to I^{2}$ be given by $h^{\pm}_{*}(a,b)=(h^{\pm}(a),h^{\pm}(b))$. If $K=(K,\ell)$ is a composite k.d., define $K_{0}=(K_{0},\ell_{0})$ and $K_{1}=(K_{1},\ell_{1})$ be the factors of $K$ defined by
$K_{0}$ $=\widehat{h}^{-}\circ K\circ(h^{-})^{-1}$ $\ell_{0}=h^{-}_{*}\circ\ell\circ(\widehat{h}^{-})^{-1}$
$K_{1}$ $=\widehat{h}^{+}\circ K\circ(h^{+})^{-1}$ $\ell_{1}=h^{+}_{*}\circ\ell\circ(\widehat{h}^{+})^{-1}.$
If $K_{0}$ and $K_{1}$ are factors of $K$, denote $K=K_{0}\mathbin{\#}K_{1}$.
Remark 2.5.
$K=K_{0}\mathbin{\#}K_{1}$ if and only if $\mu(K)=\mu(K_{0})\mathbin{\#}\mu(K_{1})$.
2.7 Crossing Number
The number of double points in a (partial) knot diagram $K$ is denoted $\chi(K)$. As usual, define
$c(K)=\min\{\chi(K^{\prime})\,:\,K^{\prime}\sim_{\mathcal{R}}K\}.$ (2.2)
For (partial) knot diagrams $P$ and $Q$ denote by $\operatorname{\mathbf{i}}(P,Q)$ the number of crossings between $P$ and $Q$ not counting $o$ or the boundary:
$\operatorname{\mathbf{i}}(P,Q)=\operatorname{Card}(P\cap Q\setminus(\partial\square\cup\{o\})).$
Here we assume that $P$ and $Q$ are mutually in general position (modulo $o$ and $\partial\square$), i.e. $\operatorname{Cross}(P)\cap\operatorname{Cross}(Q)=\varnothing$ and all intersections between $P$ and $Q$ (except for $o$ and those on the boundary) are transversal double points.
Remark 2.6.
For all knot diagrams $K$, $c(K)=c(\mu(K))$ and $\chi(K)=\chi(K^{+})+\chi(K^{-})+\operatorname{\mathbf{i}}(K^{+},K^{-})$.
3 Results
In this section we present the main results of this paper.
3.1 The Main Result
Recall the definitions of the equivalence relaions $\sim^{X}_{\mathcal{R}}$ for $X\in\{*,o1,o2,s1,s2\}$ from Section 2.5.
{restatable}
ThmMainThm Suppose that $K=K_{0}\mathbin{\#}K_{1}$ and $X\in\{*,o1,o2,s1,s2\}$. If $\widehat{K}\sim^{X}_{\mathcal{R}}K$, then $\chi(\widehat{K})\geqslant c(K_{0})+c(K_{1})$.
By Remarks 2.4, 2.5, and 2.6, the case X=X=* follows from any other case, and the cases $X=o2$ and $X=s2$ follow by symmetry from $X=o1$ and $X=s1$ respectively. So it will be enough to consider the cases $X=o1$ and $X=s1$.
3.2 Topological interpretations
The case $X=s1$ has an intuitive topological interpretation. Let $C=\square\times I=I^{3}$ be the unit cube. Let $\hat{o}^{\pm}=(o^{\pm},\tfrac{1}{2})$ be the middpoints of the upper and lower faces respectively, and let $\hat{o}=(o,\tfrac{1}{2})$ be the midpoint of the cube. Let $f^{+}=(\tfrac{1}{2},\tfrac{1}{2},1)$ be the midpoint of the frontal face of the cube. Let $W=I\times\left\{\tfrac{1}{2}\right\}\times I$ be the horizontal middle cross-section of the cube. A knot is an embedding $\mathscr{K}\colon I\to C$ such that $\mathscr{K}(0)=\hat{o}^{-}$, $\mathscr{K}(1)=\hat{o}^{+}$, $\mathscr{K}(t)\in\overset{\circ}{C}{}$ for $0<t<1$, and the one-sided derivatives at $t=0$ and $t=1$ are vertical. It is centered, if $\mathscr{K}\left(\tfrac{1}{2}\right)=\hat{o}$ and the derivative at $\tfrac{1}{2}$ is vertical. It is composite, if $\mathscr{K}\cap W=\{\hat{o}\}$. It is in general position, if $\operatorname{pr}\circ\mathscr{K}$ is a generic immersion where $\operatorname{pr}\colon C\to\square$ is the projection along the third coordinate $\operatorname{pr}(x,y,z)=(x,y)$. For a centered knot $\mathscr{K}$, denote $\mathscr{K}^{\pm}=\mathscr{K}\!\restriction\!I^{\pm}$. Let $C^{\pm}=\square^{\pm}\times I$ be the upper and lower halves of the cube. Let $g^{\pm}\colon C^{\pm}\to C$ be the linear homeomorphism given by $g(x,y,z)=(x,h(y),z)$ where $h$ is as defined in Section 2.6. If $\mathscr{K}$ is composite, let $\mathscr{K}_{0}=g^{-}\circ\mathscr{K}^{-}$ and $\mathscr{K}_{1}=g^{+}\circ\mathscr{K}^{+}$. If the knot $\mathscr{K}$ is in general position, define $\mathscr{D}(\mathscr{K})=(K,\ell)$ be the knot diagram such that $K=\operatorname{pr}\circ\mathscr{K}$ and $\ell(x)$ orders the set $K^{-1}(x)$ according to the distance from $\square\times\{0\}$. It is now clear that for a composite $\mathscr{K}$, $\mathscr{D}(\mathscr{K})$ is a composite knot diagram, and $\mathscr{D}(\mathscr{K}_{0})=\mathscr{D}(\mathscr{K})_{0}$ and $\mathscr{D}(\mathscr{K}_{1})=\mathscr{D}(\mathscr{K})_{1}$. We can now formulate a corollary to Theorem 3.1:
Corollary 1.
Suppose $\mathscr{K}$ is a centered composite knot in general position. Let $K_{0}=\mathscr{D}(\mathscr{K}_{0})$ and $K_{1}=\mathscr{D}(\mathscr{K}_{1})$. Suppose that $h\colon C\to C$ is any self-homeomorphism of $C$ which fixes pointwise the boundary $\partial C$ and the line segment $\hat{o},f^{+}]$. Let $K^{\prime}=\mathscr{D}(h\mathscr{K})$. Then $\chi(K^{\prime})\geqslant c(K_{0})+c(K_{1})$.
Proof.
It is easy to see that $\mathscr{D}(h\mathscr{K})$ can be obtained from $\mathscr{D}(\mathscr{K})=K_{0}\mathbin{\#}K_{1}$ using moves in $\mathcal{R}^{*}_{same1}$. So $K^{\prime}\sim^{s1}_{\mathcal{R}}K_{0}\mathbin{\#}K_{1}$. By Theorem 3.1, $\chi(K^{\prime})\geqslant c(K_{0})+c(K_{1})$. ∎
The cases $X=o1,o2$ also have a topological interpretation, but it is less intuitive. Let $f^{-}=(\tfrac{1}{2},\tfrac{1}{2},0)$ be the midpoint of the back face of the cube (opposite of $f^{+}$) and let $W_{d}$ be the diagonal cross section of the cube which equals to the convex hull of the set $\{(0,0,1),(1,0,1),(0,1,0),(1,1,0)\}$. We then have:
Corollary 2.
Suppose $\mathscr{K}$ is a centered composite knot in general position such that $\mathscr{K}\cap W_{d}=\{\hat{o}\}$. Let $K_{0}=\mathscr{D}(\mathscr{K}_{0})$ and $K_{1}=\mathscr{D}(\mathscr{K}_{1})$. Suppose that $h\colon C\to C$ is any self-homeomorphism of $C$ such that $h[W_{d}]\cap[f^{-},f^{+}]=\{\hat{o}\}$. Let $K^{\prime}=\mathscr{D}(h\mathscr{K})$. Then $\chi(K^{\prime})\geqslant c(K_{0})+c(K_{1})$.
Proof.
It is not hard to see that $\mathscr{D}(h\mathscr{K})$ can be obtained from $\mathscr{D}(\mathscr{K})$ using moves in $\mathcal{R}^{*}_{opp1}$. So $K^{\prime}\sim^{o1}_{\mathcal{R}}K_{0}\mathbin{\#}K_{1}$. By Theorem 3.1, $\chi(K^{\prime})\geqslant c(K_{0})+c(K_{1})$. ∎
Remark 3.1.
The result of Corollary 2 is ad hoc. Note, for example, that the set of homeomorphisms which satisfy the premise of the corollary is not closed under composition. See Question 1
3.3 Composition of Links
A version of Theorem 3.1 also holds for links in the following way. A link diagram is a triple $L=(L,M,\ell)$ where $L\colon D\to\square$ is a generic immersion, $M\in\mathbb{N}$, $D=I\cup(M\times S^{1})$ is the disjoint union of II with $M$ many circles, $L\!\restriction\!I=(L\!\restriction\!I,\ell\!\restriction\!\operatorname{Cross}(L\!\restriction\!I))$ is a knot diagram, and for all $t\in M\times S^{1}$, $L(t)\in\overset{\circ}{\square}{}$. It is centered, if $L\!\restriction\!I$ is centered. It is composite, if $L\cap H=\{o\}$. Note that it follows that $(L\!\restriction\!I)\cap H=\{o\}$ is composite (recall the definition of $H$ in the beginning of Section 2.6). A partitioned link is a tuple $(L,M^{+},M^{-},\ell)$ such that $(L,M^{+}\cup M^{-},\ell)$ is a link and $M=M^{+}\sqcup M^{-}$. For a composite $L$, there is a unique partition $M=M^{+}\sqcup M^{-}$ such that
$L\!\restriction\!(I^{\pm}\cup(M^{\pm}\times S^{1}))\subset\square^{\pm}$ (3.1)
Given a partitioned link $L$, let $L^{\pm}=(L\!\restriction\!(I^{\pm}\cup(M^{\pm}\times S^{1})))$.
Let $L=(L,M,\ell)$ be a composite link. Let $\bar{L}=(L,M^{+},M^{-},\ell)$ be the associated partitioned link with the partition witnessing (3.1), and let the homeomorphisms
$h^{\pm}_{L}\colon I^{\pm}\cup(M^{\pm}\times S1)\to I\cup(M\times S^{1})$
be defined by $h^{\pm}_{L}(t)=h^{\pm}(t)$, if $t\in I^{\pm}$ (where $h^{\pm}$ is as defined in Section 2.6) and otherwise $h^{\pm}_{L}(m,t)=(m,t)$, i.e. $h^{\pm}_{L}\!\restriction\!(M^{\pm}\times S^{1})$ is the standard inclusion into $M\times S^{1}$. Define the factors $L_{0}=(L_{0},M_{0}\ell_{0})$ and $L_{1}=(L_{1},M_{1},\ell_{1})$ of $L$ as
$L_{0}$ $=\widehat{h}^{-}\circ L\circ(h^{-}_{L})^{-1}$ $\ell_{0}=h^{-}_{*}\circ\ell\circ(\widehat{h}^{-})^{-1}$ $M_{0}=M^{-}$
$L_{1}$ $=\widehat{h}^{+}\circ K\circ(h^{+}_{L})^{-1}$ $\ell_{1}=h^{+}_{*}\circ\ell\circ(\widehat{h}^{+})^{-1}$ $M_{1}=M^{+}.$
where $\widehat{h}^{\pm}$ and $h^{\pm}_{*}$ are as in Section 2.6. If $L_{0}$ and $L_{1}$ are factors of $L$, denote $L=L_{0}\mathbin{\#}L_{1}$. The Reidemeister moves, crossing number, and minimal crossing number for links are defined just as for knots. Given a partiioned link $L$, define the restricted Reidemeister moves $R^{{}^{*}}_{\pm\mp}$ just as for knots replacing $K^{\pm}$ with $L^{\pm}$. For example in $R^{*}_{++}$ for links, the subarc of $L^{+}$ can only pass over the critical point $o$. This enables to generalize the relations $\sim^{X}_{\mathcal{R}}$, $X\in\{*,o1,o2,s1,s2\}$ for links. We can now state a version of Theorem 3.1 for links:
{restatable}
ThmMainThmLinks Suppose that $L=L_{0}\mathbin{\#}L_{1}$ and $X\in\{*,o1,o2,s1,s2\}$. If $L^{\prime}\sim^{X}_{\mathcal{R}}\bar{L}$, then $\chi(L^{\prime})\geqslant c(L_{0})+c(L_{1})$.
We will dedicate this paper to proving Theorem 3.1 and explain in Section 9 how to adapt the proof to links.
Corollary 3 (Topological interpretation for links).
Corollaries 1 and 2 hold for link compositions instead of knot compositions.
3.4 A Version for Knots Diagrams on $S^{2}$
The results and proofs in this paper are stated for knots presented as tangle diagrams, but they apply for standard knots through a standard transformation. Parametrize the unit circle as $S^{1}=\{e^{i\theta}\,:\,0\leqslant\theta\leqslant 2\pi\}$. Denoet the upper and lower halves of the circle by $S^{1}_{+}=\{e^{i\theta}\,:\,0\leqslant\theta\leqslant\pi\}$ and $S^{1}_{-}=\{e^{i\theta}\,:\,\pi\leqslant\theta\leqslant 2\pi\}$. Think of $S^{2}=\mathbb{R}^{2}\cup\{\infty\}$. Let $\mathbb{R}_{\geqslant 0}=\{x\in\mathbb{R}\,:\,x\geqslant 0\}$ and $\mathbb{R}_{\leqslant 0}=\{x\in\mathbb{R}\,:\,x\leqslant 0\}$. A (classical) knot diagram is a generic embedding $K\colon S^{1}\to S^{2}$ equipped with the under and over crossing information. W.l.o.g. assume that $\infty\notin|K|$. The notions $c(K)$ and $\chi(K)$ are defined in the same way as previously. Suppose that $K$ satisfies the following:
$K(t)\in\mathbb{R}\times\{0\}\iff t\in\{e^{i0},e^{i\pi}\},\quad K(e^{i0})=(0,0),\quad\text{ and }\quad\ K(e^{i\pi})=(1,0).$
Then the diagram is called composite. In this case, let
$K_{0}$ $=K\cap(R\times\mathbb{R}_{\leqslant 0})\cup([0,1]\times\{0\}),\ \text{and}$
$K_{1}$ $=K\cap(R\times\mathbb{R}_{\geqslant 0})\cup([0,1]\times\{0\}).$
After reparametrization (cf. Section 4.5) these are factors of $K$, and we denote $K=K_{0}\mathbin{\#}K_{1}$. Define restricted Reidemeister moves for such diagrams in the same way as in Section 2.5 with the difference that instead of one critical point $o$ we have two critical points, $(0,0)$ and $(1,0)$. To be more formal, let $\mathcal{R}^{c}$ be the set of those (classical) Reidemeister moves which are performed in a neighborhood which does not contain either of the critical points. A sliding move is performed in a neighborhood which contains exactly one of the critical points and in such a move an arc slides over or under that point. Let $R^{c}_{\pm\mp}$ be the sliding move where a subarc of $K\!\restriction\!S^{1}_{\pm}$ is only allowed to cross over (if $\mp=+$) or under (if $\mp=$-) a critical point. Then define $\mathcal{R}^{c}_{opp1}$, $\mathcal{R}^{c}_{opp2}$, $\mathcal{R}^{c}_{same1}$, and $\mathcal{R}^{c}_{same2}$ by analogy with $\mathcal{R}^{*}_{opp1}$, $\mathcal{R}^{*}_{opp2}$ etc. This induces the relations $\sim^{o1}_{\mathcal{R}}$, $\sim^{o2}_{\mathcal{R}}$, $\sim^{s1}_{\mathcal{R}}$, and $\sim^{s2}_{\mathcal{R}}$ for classical knot diagrams. We can then formulate:
Corollary 4.
Suppose $K=K_{0}\mathbin{\#}K_{1}$ is a composite (classical) knot diagram on $S^{2}$ and $X\in\{*,o1,o2,s1,s2\}$. If $\widehat{K}\sim^{X}_{\mathcal{R}}K$, then $\chi(\widehat{K})\geqslant c(K_{0})+c(K_{1})$.
Proof.
By a standard argument one can assume that the point $(1,0)$ is never crossed, i.e. all sliding moves over/under that point can be replaced by a sequence of other moves. Instead of crossing $(1,0)$ one can move the arc over the other side of $S^{2}$, passing through $\infty$. W.l.o.g. one can assume that a neighborhood of $\infty$ is never occupied by $K$ (before or after a restricted Reidemeister move). Cutting open at that neighborhood and projecting onto $\square$, one obtains knot diagram as in the previous sections (as well as the rest of the paper). Crossing the “other” critical point becomes a move as shown on Figure 3. This translates the problem back to the one presented above and so we can apply Theorem 3.1. ∎
Figure 3: Crossing the infinity point on $S^{2}$ translated to a tangle diagram.
[…]
4 Basics of Tangle Diagrams
We will need to deal with tangle diagrams which are generalizations of knot and link diagrams defined in Section 3.
4.1 Definition of Tangle Diagrams
A tangle diagram (a t-diagram for short) is a tuple $T=(T,N,M,\ell)$ such that
  1. ($T_{1}$)
    NN and $M$ are disjoint finite sets.
  2. ($T_{2}$)
    $T$ is a generic immersion $T\colon D_{T}\to\square$ where $D_{T}=(I\times N)\cup(S^{1}\times M)$ is the disjoint union of $\operatorname{Card}(N)$ many unit intervals and $\operatorname{Card}(M)$ many circles.
  3. ($T_{3}$)
    $|T|$ is disjoint from the corners of the square, $|T|\cap\{(0,0),(0,1),(1,0),(1,1)\}=\varnothing$.
  4. ($T_{4}$)
    $T^{-1}[\partial\square]\subset\{0,1\}\times N$, i.e. a point is mapped into the boundary $\partial\square$ only if it is an endpoint of one of the strands, or equivalently a boundary point of the manifold $D_{T}$.
  5. ($T_{5}$)
    $\ell\colon\operatorname{Cross}(T)\to D_{T}^{2}$ is a labeling function which tells which arc is above the other in a crossing. For each $x\in\operatorname{Cross}(T)$, we have $\ell(x)=(a,b)$ where $a$ and $b$ are such that $\{a,b\}=T^{-1}(x)$
Note that it follows from the definition of transversality that $\operatorname{Cross}(T)\cap\partial\square=\varnothing$. As for knot and link diagrams, the number of crossings in $T$ is denoted $\chi(T)=\operatorname{Card}(\operatorname{Cross}(T))$. The restrictions $T\!\restriction\!(I\times\{i\})$ for $i\in N$ are the strands of $T$ and the restrictions $T\!\restriction\!(S^{1}\times\{j\})$ for $j\in M$ are the loops. If $T=(T,N,M,\ell)$ is a t-diagram and $S=T\!\restriction\!(I\times\{i\})$ is a strand of $T$, then $(S,\{i\},\varnothing,\ell_{S})$ is also a t-diagram where $\ell_{S}=\ell_{T}\!\restriction\!\operatorname{Cross}(S)$. Similarly, if $j\in M$, $(S,\varnothing,\{j\},\ell_{S})$ is a t-diagram where $S=T\!\restriction\!(S^{1}\times\{j\})$. Denote by $E(T)=|T|\cap\partial\square$ the set of endpoints. A t-diagram is loopless if $M=\varnothing$. In this case we may drop $M$ from the notation and consider triples $(T,N,\ell)$. A knot diagram is a special case of a loopless t-diagram $T$ with $N=1$, and a link diagram is a special case of a (not-necessarily-loopless) t-diagram with $N=1$. As mentioned in Section 2.1, we often identify $T$ with its range $|T|$ and drop the vertical bars from the notation.
Note that in ($T_{4}$) we did not require equivalence (we required “only if” as opposed to “if and only if”), so there can be loose ends in a t-diagram, i.e. boundary points of $D_{T}$ which are mapped into the interior of $\square$. Thus, the partial diagrams defined in Section 2.4 are also special cases of t-diagrams. A t-diagram is called complete if there are no loose ends and $T^{-1}[\partial\square]=\partial D_{T}$. For a complete t-diagram we have $\operatorname{Card}(E(T))=2\operatorname{Card}(N)$.
4.2 Operations on Tangle Diagrams
Definition 4.1.
A rectangle $A\subset\mathbb{R}^{2}$ is axis-aligned if it is of the form $a,b]\times[c,d]$ for some $a<b$ and $c<d$. A homeomorphism $h\colon A\to B$ between axis-aligned rectangles $A=[a,b]\times[c,d]$ and $A^{\prime}=[a^{\prime},b^{\prime}]\times[c^{\prime},d^{\prime}]$ is called axis-aligned if there are increasing homeomorphisms $h_{1}\colon[a,b]\to[a^{\prime},b^{\prime}]$ and $h_{2}\colon[c,d]\to[c^{\prime},d^{\prime}]$ such that $h(x,y)=(h_{1}(x),h_{2}(y))$ for all $(x,y)\in A$.
Fact 4.2.
Given two axis-aligned rectangles $A,A^{\prime}$ there is a unique linear axis-aligned homeomorphism between them. ∎
Definition 4.3.
Let $G$ be an axis-aligned rectangle. A tuple $(T,N,M,\ell)$ is a $G$-tangle diagram ($G$-t-diagram for short) if $T\subset G$ and $(h\circ T,N,M,\ell\circ h^{-1})$ is a t-diagram where $h$ is the unique linear axis-aligned homeomorphism $h\colon G\to\square$.
Definition $4.$
Suppose $T=(T,N,M,\ell)$ is a t-diagram and $G\subset\square$ an axis-aligned rectangle. We say that $T$ is in general position with respect to $G$ if $\operatorname{Cross}(T)\cap\partial G=\varnothing$, $T$ does not intersect $G$ in the corners, and each intersection of $T$ with $\partial G$ is transversal. In this case, $T^{-1}[G]$ is a finite union of nondegenerate connected closed submanifolds of $D_{T}$. Thus, there exist $N^{\prime}$ and $M^{\prime}$, and a homeomorphism $\eta\colon(I\times N^{\prime})\cup(S^{1}\times M^{\prime})\to D_{T}$. Define
$T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$
to be the $G$-t-diagram $(T\circ\eta,N^{\prime},M^{\prime},\eta^{-1}\circ(\ell\!\restriction\!G))$. Note that the choice of $\eta$ is not unique, but it does not matter for our purposes. The choice of $N^{\prime}$ and $M^{\prime}$ is unique.
4.3 Type of a Tangle
Denote the vertical and horizontal open edges of the square as follows:
$V^{-}$ $=\{(0,t)\,:\,0<t<1\}$ $=\text{vertical left edge}$
$V^{+}$ $=\{(1,t)\,:\,0<t<1\}$ $=\text{vertical right edge}$
$H^{-}$ $=\{(t,0)\,:\,0<t<1\}$ $=\text{horizontal lower edge}$
$H^{+}$ $=\{(t,1)\,:\,0<t<1\}$ $=\text{horizontal upper edge}$
Note that the square’s corner points are not included in these sets.
Definition 4.5.
Let $n_{1},n_{2},n_{3},n_{4}\in\mathbb{N}$. We say that a t-diagram $T=(T,N,M,\ell)$ is of type $(n_{1},n_{2},n_{3},n_{4})$ if
$\operatorname{Card}(T\cap H^{+})=n_{1},\ \operatorname{Card}(T\cap V^{-})=n_{2},\ \operatorname{Card}(T\cap H^{-})=n_{3},\text{ and }\operatorname{Card}(T\cap V^{+})=n_{4}.$
Recall that the t-diagram avoids the corners and there are no multiplicity points on $\partial\square$, so we necessarily have $n_{1}+n_{2}+n_{3}+n_{4}=\operatorname{Card}(E(T))$, and if $T$ is complete, then also $2\operatorname{Card}(N)$.
Definition 4.6.
Suppose $T=(T,N,M,\ell)$ is a t-diagram and let $E_{V}(T)=T\cap(V^{-}\cup V^{+})$ be the set of all its endpoints which are on the vertical edges of $\square$. The set $(V^{-}\cup V^{+})\setminus E_{V}(T)$ is the union of disjoint open subarcs of $V^{-}\cup V^{+}$. Let $L_{1}(T),\dots,L_{m}(T)$ be the counterclockwise enumeration of these arcs so that $L_{1}(T)$ is the one whose closure contains the left top corner $(0,1)$. See Figure 4 for illustration.
Figure 4: The sets $L_{1}(T),\dots,L_{m}(T)$ of a given t-diagram illustrated. They list the connected components of the vertical edges of the square from which the endpoints of the tangle have been removed. The enumeration is counterclockwise starting from the top left corner.
Definition 4.7.
A special case of this was already considered in Section 2.7. Two t-diagrams $T$ and $S$ are in mutual general position if $E(S)\cap E(T)=\varnothing$, $T\cap\operatorname{Cross}(S)=S\cap\operatorname{Cross}(T)=\varnothing$, there are no more than finitely many $x\in T\cap S$ and each such $x$ is a transversal double point between an arc of $T$ and an arc of $S$ unless $x$ is a loose point in which case we require that it is a loose point of both. For example, if $K$ is a knot diagram (Section 2.3), then $K^{+}$ and $K^{-}$ are in mutual general position, even though $o\in K^{+}\cap K^{-}$ is not a transversal double point. When two t-diagrams $T$ and $S$ are in mutual general position, denote
$\operatorname{\mathbf{i}}(T,S)=\operatorname{Card}(\overset{\circ}{T}{}\cap\overset{\circ}{S}{})$ (4.1)
where by $\overset{\circ}{T}{}$ in this context we mean the restriction of $T$ to $D_{T}\setminus\partial D_{T}=$ the interior points of the manifold precisely because we do not want $\operatorname{\mathbf{i}}(\cdot,\cdot)$ to count the possible loose points. We always assume by default that all t-diagrams are in mutual general position in a given context.
4.4 Applying a Homeomorphism to a Tangle
Suppose $h\colon\square\to\square$ is a homeomorphism and $(T,N,M,\ell)$ a t-diagram. Then we denote by hThT the t-diagram $(h\circ T,N,M,\ell\circ h^{-1})$. Important special cases are the applications of the moves $R_{0}$ and $R^{*}_{0}$, (Section 2.5).
4.5 Reparametrization
Sometimes t-diagrams need to be reparametrized. An example of this was the restriction of a t-diagram to a rectangle $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$, Definition $4.$ Another example is smoothing. A local smoothing of a t-diagram is a local change at a crossing as depicted in Figure 6. If t-diagrams $(T,N,M,\ell)$ and $(T^{\prime},N^{\prime},M^{\prime},\ell^{\prime})$ are the same t-diagram before and after smoothing, the number of strands and loops may or may not change, i.e. we might have $N\neq N^{\prime}$ and/or $M\neq M^{\prime}$, or they might not change etc. We will not give the technical details of the reparametrization and assume that the reader understands how it can be done in each specific case. A smoothing operation may connect two strands which were originally oriented in opposite directions. Then one of them changes orientation after merging with the other strand. This is not a problem, as in this context we do not care about the orientation of strands and loops in t-diagrams.
Similarly for curves. We introduced the notations $\gamma_{1}{}^{\frown}\gamma_{2}$ and $f\cup g$ in the beginning of this section, but sometimes the domain of the curves may not be exactly equal to II, or they may be oriented in the “wrong” direction, i.e. $\gamma_{1}(0)=\gamma_{2}(0)$ and we want to connect them at their beginning points. It should always be clear from the context how to (re-)parametrize the curves in a proper manner.
In above examples the orientation is not of crucial importance. Sometimes, however, we use orientation and related properties as a technical tool. For example in the definition of a centered knot diagram (see Sections 2.3 and 2.4) we required that the points $0,\frac{1}{2},1\in I$ are always mapped respectively to the bottom, the middle, and the top of the square which indirectly implies an orientation.
5 Bounding the Number of Non-Hybrid Crossings
To prove Theorem 3.1, we will first bound the number of the “non-hybrid crossings”, i.e. the crossings in $\widehat{K}^{+}$ and $\widehat{K}^{-}$. It is done in this section and is the easier part of the proof of Theorem 3.1.
5.1 The Minimal Ray Intersection Number
Definition 5.1.
A ray is curve without self-intersections in $\square$ which has one endpoint at $o$ and another at $\partial\square$. For a t-diagram $T$, let
$\operatorname{MIN}_{\operatorname{ray}}(T)$ $=\min\{\operatorname{\mathbf{i}}(P,J)\,:\,J\text{ is a ray}\}.$
Call it the minimal ray intersection number. As usual, we only consider rays which are in general position with respect to $T$. Note that by the definition of $\operatorname{\mathbf{i}}(\cdot,\cdot)$, (4.1), the point $o$ is not counted in $\operatorname{\mathbf{i}}(T,J)$.
Recall that by $\sim^{*}_{0}$ we denote the equivalance relation on (partial) knot diagrams which is induced by the $R_{0}$-moves alone, i.e. $K\sim^{*}_{0}K^{\prime}$ iff there is a self-homeomorphism $h\colon\square\to\square$ such that $h$ is the identity on some neighborhood of $\partial\square\cup\{o\}$ and $K^{\prime}=hK$.
Lemma 5.2 (Invariance under $R_{0}$).
Let $K$ be a centered knot diagram and $\widehat{K}\sim^{*}_{0}K$. Then $\operatorname{MIN}_{\operatorname{ray}}(\widehat{K}^{\pm})=\operatorname{MIN}_{\operatorname{ray}}(K^{\pm})$.
Proof.
A sequence of $R_{0}$-moves is equivalent to one $R_{0}$-move, because the composition of witnessing homeomorphisms is a witnessing homeomorphism. So suppose that $\widehat{K}$ has been obtained from $K$ using such a homeomorphism $h$. Recall that by the definition of the $R_{0}$-move, $h$ keeps $o$ fixed, $J$ is a ray if and only if hJhJ is a ray and $\operatorname{\mathbf{i}}(K^{\pm},J)=\operatorname{\mathbf{i}}(hK^{\pm},hJ)$. So the statement follows from the obvious equality $hK^{\pm}=\widehat{K}^{\pm}$. ∎
5.2 The Bound on $\chi(K^{\pm})$
Recall the definition of a composite knot diagram from Section 2.6 and the definition of the equivalence relations $\sim^{X}_{\mathcal{R}}$ for $X\in\{*,o1,o2,s1,s2\}$ from Section 2.5. Recall also the notations $\hat{o}$ and $\hat{o}^{\pm}$ from Section 3.2.
Proposition 5.1.
Suppose $K$ is a composite centered knot diagram and suppose $\widehat{K}\sim^{X}_{\mathcal{R}}K$ for some $X\in\{*,o1,o2,s1,s2\}$. Let $K_{i}$, $i\in\{0,1\}$, be the factors of $K$, i.e. $K=K_{0}\mathbin{\#}K_{1}$. Then there are knot diagrams $\widehat{K}_{0}$ and $\widehat{K}_{1}$ such that $\widehat{K}_{0}\sim_{\mathcal{R}}K_{0}$, $\widehat{K}_{1}\sim_{\mathcal{R}}K_{1}$, and the following equalities hold:
$\chi(\widehat{K}_{1})$ $=\chi(\widehat{K}^{+})+\operatorname{MIN}_{\operatorname{ray}}(\widehat{K}^{+})$ (5.1)
$\chi(\widehat{K}_{0})$ $=\chi(\widehat{K}^{-})+\operatorname{MIN}_{\operatorname{ray}}(\widehat{K}^{-})$ (5.2)
Proof.
The case X=X=* will follow from any of the other cases. Suppose that $X=o1$. Then the subarcs of $K^{+}$ are not allowed to slide under the critical point $o$. Let $\mathscr{K}$ and $\widehat{\mathscr{K}}$ be knots in $C=I^{3}$ such that $\mathscr{D}(\mathscr{K})=K$ and $\mathscr{D}(\widehat{\mathscr{K}})=\widehat{K}$. Then there is a boundary-fixing homeomorphism $h\colon C\to C$ such that $\widehat{\mathscr{K}}=h\widehat{\mathscr{K}}$. By the assumption, we can assume that $h$ fixes the tube $\tau=B(o,\varepsilon)\times I^{-}$ for some sufficiently small $\varepsilon$. We can also assume that there is a neighborhood $U$ of $\partial C$ which is fixed pointwise by $h$. Let $J$ be a ray such that $\operatorname{\mathbf{i}}(K^{+},J)=\operatorname{MIN}_{\operatorname{ray}}(K^{+})$. By modifying $J$ in a neighborhood of $\partial\square$, we can assume without loss of generality that the other endpoint of $J$ is at $o^{-}$ and that $(J,\ell_{\varnothing})$ satisfies the conditions of being a lower partial knot diagram without crossings. It is now easy to find a curve $\mathscr{J}$ in $C$ such that one endpoint of $\mathscr{J}$ is at $\hat{o}$, the other endpoint is on $\partial C$, $\mathscr{J}$ is contained in the set $\tau\cup U$, and $\operatorname{pr}\mathscr{J}=J$. Let $\mathscr{K}_{1}=\mathscr{K}^{+}\cup\mathscr{J}$ and $\widehat{\mathscr{K}}_{1}=\widehat{\mathscr{K}}^{+}\cup\mathscr{J}$. We now have $\widehat{K}K_{1}=h\mathscr{K}_{1}$. On the other hand, it is also clear that $\mathscr{D}(\mathscr{K}_{1})$ is equivalent to the composition of $K_{1}$ with a trivial knot, so $\mathscr{D}(\mathscr{K}_{1})\sim_{\mathcal{R}}K_{1}$ and hence also $\mathscr{D}(\widehat{\mathscr{K}}_{1})\sim_{\mathcal{R}}K_{1}$. Let $\widehat{K}_{1}=\mathscr{D}(\widehat{\mathscr{K}}_{1})$. So we have $\widehat{K}_{1}\sim_{\mathcal{R}}K_{1}$. To compute the number of crosings in $\widehat{K}_{1}$, note that $\widehat{K}_{1}^{-}=J$, so by Remark 2.6,
$\chi(\widehat{K}_{1})=\chi(\widehat{K}_{1}^{+})+\chi(J)+\operatorname{\mathbf{i}}(\widehat{K}^{+}_{1},J)$
But $\widehat{K}_{1}^{+}=\widehat{K}^{+}$, $\chi(J)=0$, and $\operatorname{\mathbf{i}}(\widehat{K}^{+},J)=\operatorname{MIN}_{\operatorname{ray}}(\widehat{K}^{+})$, so we have
$\chi(\widehat{K}_{1})=\chi(\widehat{K}^{+})+\operatorname{MIN}_{\operatorname{ray}}(\widehat{K}^{+})$
as required. By a symmetric argument, obtain $\widehat{K}_{0}$ with $\widehat{K}_{0}\sim_{\mathcal{R}}K_{0}$ which satisfies (5.2). The cases $X\in\{o2,s1,s2\}$ are obtained by symmetric arguments. ∎
Corollary 5 (A Bound for Non-Hybrid Crossings).
Suppose that $K$ is a composite centered knot diagram and $K=K_{0}\mathbin{\#}K_{1}$. Suppose that $\widehat{K}\sim^{X}_{\mathcal{R}}K$ for some $X\in\{*,o1,o2,s1,s2\}$. Then
$\chi(\widehat{K}^{+})+\operatorname{MIN}_{\operatorname{ray}}(\widehat{K}^{+})$ $\geqslant c(K_{1})$ (5.3)
$\chi(\widehat{K}^{-})+\operatorname{MIN}_{\operatorname{ray}}(\widehat{K}^{-})$ $\geqslant c(K_{0})$ (5.4)
Proof.
Let $\widehat{K}_{0}$ and $\widehat{K}_{1}$ be as given by Proposition 5.1. Since $\widehat{K}_{0}\sim_{\mathcal{R}}K_{0}$ and $\widehat{K}_{0}\sim_{\mathcal{R}}K_{0}$, it follows (from (2.2)) that $\chi(\widehat{K}_{0})\geqslant c(K_{0})$ and $\chi(\widehat{K}_{1})\geqslant c(K_{1})$. The statement follows now from (5.1) and (5.2). ∎
6 Crossings Between Overlapping Tangle Diagrams
We have now established a lower bound on the number of non-hybrid crossings in $\hat{K}$. Next, we will have to establish a lower bound for the number of hybrid crossings $\operatorname{\mathbf{i}}(\widehat{K}^{-},\widehat{K}^{+})$. Together they will lead to the proof of Theorem 3.1 using the equation in Remark 2.6. This will be done in Section 7. In this section we prove preliminary results needed. The main objective is to prove Proposition 6.1 which gives a lower bound for the number of crossings between two “overlapping” tangle diagrams under certain conditions. To prove it, we will need a lemma about integer-valued matrices (Section 6.1, Lemma 6.5) and a lemma about tangle diagrams which enables the application of Lemma 6.5 to the context of t-diagrams (Section 6.2, Lemma 6.9).
6.1 Multigraphs, Matrices and Zigzag Paths
Denote by $\mathbb{N}$ natural numbers including zero and by $\mathbb{N}_{+}=\mathbb{N}\setminus\{0\}$ natural numbers excluding zero. For $n,m\in\mathbb{N}$ denote by $\llbracket n,m\rrbracket$ the discrete set $\{n,n+1,\dots,m-1,m\}$. By convention, if $m<n$, then $\llbracket n,m\rrbracket$ is the empty set and if $m=n$, then it is the singleton $\{n\}=\{m\}$. Recall that $\operatorname{Card}(A)$ the cardinality of the set AA, e.g. for $n\leqslant m$ we have
$\operatorname{Card}(\llbracket n,m\rrbracket)=m-n+1.$ (6.1)
The ijij-th element of an $(n\times n)$-matrix $e$ is expressed as $e(i,j)$ where $i,j\in\llbracket 1,n\rrbracket$. Let $\mathcal{M}^{s}_{n}(\mathbb{N})$ be the set of symmetric $\mathbb{N}$-valued $(n\times n)$-matrices,
$\mathcal{M}^{s}_{n}(\mathbb{N})=\{x\in\mathbb{N}^{(n\times n)}\,:\,\forall i,j\in\llbracket 1,n\rrbracket(x(i,j)=x(j,i))\}.$
A useful intuitive interpretation of a square $\mathbb{N}$-valued matrix is that it represents a multigraph. In this interpretation, the value $e(i,j)$ is the number of connections from node $i$ to node $j$. This is, intuitively, how the results of this section are going to be applied in the context of t-diagrams (see Definition 6.22).
Throughout this section we fix $k\in\mathbb{N}_{+}$ and let $n=2k$. Let $f\colon\llbracket 1,n\rrbracket\to\llbracket 1,n\rrbracket$ be the function defined by f(i)=ni+1f(i)=n-i+1. Note that
$f(f(i))=i$ (6.2)
for all $i\in\llbracket 1,n\rrbracket$, in particular $f$ is a bijection. We will also need the fact that
$i\in\llbracket 2,k-1\rrbracket\iff f(i)\in\llbracket k+2,n-1\rrbracket.$ (6.3)
For an $(n\times n)$-matrix $e$, denote by $s_{i}(e)$ the sum of the $i$-th row of $e$,
$s_{i}(e)=\sum_{j=1}^{n}e(i,j).$ (6.4)
In the multigraph interpretation, $s_{i}(e)$ is the total number of out-going connections from ii.
Definition 6.1.
A matrix $e\in\mathcal{M}^{s}_{n}(\mathbb{N})$ is said to satisfy property ()(*) if the following conditions hold:
  1. ($enumi_{1}$)
    For all $i\in\llbracket 2,k-1\rrbracket$, $s_{i}(e)=s_{f(i)}(e)$
  2. ($enumi_{2}$)
    If $k>1$, then $s_{1}(e)=s_{f(1)}(e)+1$ and $s_{k}(e)=s_{f(k)}(e)+1$
  3. ($enumi_{3}$)
    For all $i\in\llbracket 1,n\rrbracket$, $e(i,i)$ is even.
Definition 6.2.
Given $e\in\mathcal{M}^{s}_{n}(\mathbb{N})$, a zigzag $e$-path is a sequence $(i_{1},\dots,i_{p})\subset\llbracket 1,n\rrbracket$ such that for all $q\in\llbracket 1,p-1\rrbracket$, $e(f(i_{q}),i_{q+1})>0$.
For $a,b\in\mathcal{M}^{s}_{n}(\mathbb{N})$, denote $a\leqslant b$ if $a(i,j)\leqslant b(i,j)$ holds for all $i,j\in\llbracket 1,n\rrbracket$, and $a<b$ if $a\leqslant b$ and $a\neq b$.
Lemma 6.3.
Suppose that $e\in\mathcal{M}^{s}_{n}(\mathbb{N})$ satisfies property ()(*). Then there is a zigzag $e$-path $(i_{1},\dots,i_{p})$ such that $i_{1}=f(1)$ and $i_{p}\in\{k,f(k)\}$.
Proof.
In case $k=1$, let $P=(i_{1})$ be the path of length $p=1$, in which $i_{1}=f(1)=n=2$. Then $P$ is trivially a zigzag $e$-path, and $i_{1}=f(1)=f(k)=i_{p}$, so we can assume that $k>1$. Suppose for a contradiction that there is no zigzag path as required. We will inductively define an infinite sequence $(x_{i})_{i\in\mathbb{N}}\subset\mathcal{M}^{s}_{n}(\mathbb{N})$ such that $x_{1}>x_{2}>\cdots$. This is a contradiction, because there cannot be infinite strictly descending sequences in $\mathcal{M}^{s}_{n}(\mathbb{N})$ due to the well-foundedness of $(\mathbb{N},<)$. If $m\in\llbracket 1,n\rrbracket\setminus\{k,k+1\}$, we say that $x\in\mathcal{M}^{s}_{n}(\mathbb{N})$ satisfies the $m$-modified property ()(*) called $(*^{m})$ if:
  1. ($enumi^{m}_{1}$)
    For all $i\in\llbracket 1,n\rrbracket\setminus\{m,f(m),k,k+1\}$, $s_{i}(x)=s_{f(i)}(x_{q})$
  2. ($enumi^{m}_{2}$)
    $s_{f(m)}(x)=s_{m}(x)+1$ and $s_{k}(x)=s_{f(k)}(x)+1$.
  3. ($enumi^{m}_{3}$)
    (same as ($enumi_{3}$)): For all $i\in\llbracket 1,n\rrbracket$, $x(i,i)$ is even.
Let $P_{1}=(i_{1})$ and $X_{1}=(x_{1})$, in which $i_{1}=f(1)=n$ and $x_{1}=e$. Now using (6.2) and (6.3), it is easy to verify that $(*^{i_{1}})=(*^{n})$ is the same as ()(*), so it applies to $e$. It is also clear that $P_{1}$ is a zigzag $e$-path. Suppose that we have defined $P_{q}=(i_{1},\dots,i_{q})$ and $X_{q}=(x_{1},\dots,x_{q})$ such that $P$ is a zigzag $e$-path, $x_{1}>x_{2}>\dots$, and $x_{q}$ satisfies $(*^{m})$ for $m=i_{q}$ (note that by the counterassumptions, $i_{q}\notin\{k,k+1\}$). Then by ($enumi^{m}_{2}$) for $x_{q}$ we have $s_{f(m)}(x_{q})=s_{m}(x_{q})+1$. Since $s_{m}(x_{q})$ is non-negative, $s_{f(m)}(x_{q})>0$. By (6.4) this means that there must exist $m^{\prime}\in\llbracket 1,n\rrbracket$ with
$x_{q}(f(m),m^{\prime})>0.$ (6.5)
Let $i_{q+1}=m^{\prime}$. Since $e=x_{1}>x_{q}$, also $e(f(m),m^{\prime})>0$ holds, so $P_{q+1}=(i_{1},\dots,i_{q},i_{q+1})$ is a zigzag $e$-path. Note that by our counterassumption,
$m^{\prime}\notin\{k,f(k)\}.$ (6.6)
Let $x_{q+1}$ be defined as follows
$x_{q+1}(i,j)=$ $x_{q}(i,j)-1$ if $(i,j)\in\{(f(m),m^{\prime}),(m^{\prime},f(m))\}$ and $f(m)\neq m^{\prime}$,
$x_{q+1}(i,j)=$ $x_{q}(i,j)-2$ if $(i,j)\in\{(f(m),m^{\prime}),(m^{\prime},f(m))\}$ and $f(m)=m^{\prime}$,
$x_{q+1}(i,j)=$ $x_{q}(i,j)$ otherwise.
Clearly $x_{q}>x_{q+1}$. By (6.5), ($enumi^{m}_{3}$), and the symmetry of $x_{q}$ we have $x_{q+1}\in\mathcal{M}^{s}_{n}(\mathbb{N})$. It remains to verify that $x_{q+1}$ satisfies $(*^{m^{\prime}})$. The condition $(*^{m^{\prime}}_{3})$ for $x_{q+1}$ follows trivially from the assumption that ($enumi^{m}_{3}$) holds for $x_{q}$. So let us prove $(*^{m^{\prime}}_{1})$ and $(*^{m^{\prime}}_{2})$ for $x_{q+1}$. Suppose first that $f(m)\neq m^{\prime}$. Then the definition of $x_{q+1}$ means that the number 11 has been subtracted from exactly two rows in $x_{q}$ to obtain $x_{q+1}$. These rows are $f(m)$ and $m^{\prime}$. This means that
$s_{f(m)}(x_{q+1})$ $=s_{f(m)}(x_{q})-1$ (6.7)
$s_{m^{\prime}}(x_{q+1})$ $=s_{m^{\prime}}(x_{q})-1$ (6.8)
$s_{i}(x_{q+1})$ $=s_{i}(x_{q})$ $\text{for all }i\notin\{f(m),m^{\prime}\}$ (6.9)
By the induction hypothesis we have
$s_{f(m)}(x_{q})$ $=s_{m}(x_{q})+1,$ (6.10)
$s_{k}(x_{q})$ $=s_{f(k)}(x_{q})+1$ (6.11)
$s_{i}(x_{q})$ $=s_{f(i)}(x_{q})$ $\text{for all }i\notin\{m,f(m),k,f(k)\}.$ (6.12)
To verify $(*^{m^{\prime}}_{1})$, suppose that
$i\in\llbracket 1,n\rrbracket\setminus\{m^{\prime},f(m^{\prime}),k,f(k)\}$ (6.13)
If $i\notin\{m,f(m)\}$, then
[…]
Condition (6.9) can be applied in the last equality, because $i\notin\{f(j),j\}$ implies $f(i)\notin\{f(j),j\}$ for any $j$ by (6.2). Suppose $i=m$. Note that then $m\neq m^{\prime}$ by the choice of $i$ (6.19). Also then $i\neq f(m)$, so we have
[…]
Suppose that $i=f(m)$. By the choice of $i$ this implies again $m\neq m^{\prime}$, because $i$ cannot be equal to $f(m^{\prime})$. Note that by (6.2) we have now $f(i)=m$, so:
[…]
This proves $(*^{m^{\prime}}_{1})$ for $x_{q+1}$. Now consider $(*^{m^{\prime}}_{2})$. We are still working under the assumption that $m^{\prime}\neq f(m)$. Since $m,m^{\prime}\notin\{k,f(k)\}$, it also follows that $f(m)\notin\{k,f(k)\}$ and symmetrically that $k\notin\{m,f(m),m^{\prime},f(m^{\prime})\}$, so
[…]
This verifies the second part of ($enumi^{m}_{2}$). For the first part, first consider the case $m\neq m^{\prime}$. Then $f(m)\neq f(m^{\prime})$ and by (6.6) and the assumption $m^{\prime}\neq f(m)$ we have
[…]
Suppose that $m=m^{\prime}$. Then $f(m)=f(m^{\prime})$ and we have
[…]
We have proved $(*^{m^{\prime}})$ for $x_{q+1}$ in the case $m^{\prime}\neq f(m)$. Let us now assume that $f(m)=m^{\prime}$. Then the definition of $x_{q+1}$ means that the number 22 has been subtracted frmo exactly one row of $x_{q}$ to obtain $x_{q+1}$ that row being $m^{\prime}=f(m)$. This means:
$s_{m^{\prime}}(x_{q+1})$ $=s_{m^{\prime}}(x_{q})-2=s_{f(m)}(x_{q})-2$ (6.14)
$s_{i}(x_{q+1})$ $=s_{i}(x_{q})$ $\text{for all }i\neq m^{\prime}.$ (6.15)
Substituting $m^{\prime}$ for $f(m)$ in (6.10)–(6.12) the induction hypothesis gives:
$s_{m^{\prime}}(x_{q})$ $=s_{m}(x_{q})+1,$ (6.16)
$s_{k}(x_{q})$ $=s_{f(k)}(x_{q})+1$ (6.17)
$s_{i}(x_{q})$ $=s_{f(i)}(x_{q})$ $\text{for all }i\notin\{m,m^{\prime},k,f(k)\}.$ (6.18)
To verify $(*^{m^{\prime}}_{1})$, suppose that
$i\in\llbracket 1,n\rrbracket\setminus\{m^{\prime},f(m^{\prime}),k,f(k)\}.$ (6.19)
The assumption $m^{\prime}=f(m)$ imples $f(m^{\prime})=m$, so we also have $i\neq m$. By (6.18) we have $s_{i}(x_{q})=s_{f(i)}(x_{q})$. Since $i\notin\{m^{\prime},f(m^{\prime})\}$, also $f(i)\notin\{m^{\prime},f(m^{\prime})\}$, so (6.15) implies the desired equality. For $(*^{m^{\prime}}_{2})$, consider:
[…]
The part for $k$ follows as above because again $m^{\prime}=f(m)\notin\{k,f(k)\}$. ∎
Definition 6.4.
Let $w\in\mathcal{M}^{s}_{n}(\mathbb{N})$ be defined for all $i\leqslant j\in\llbracket 1,n\rrbracket$ by
$w(i,j)=$ $|j-i|$ if $i\leqslant j\leqslant k$ or $k<i\leqslant j$ (6.20)
$w(i,j)=$ $|f(j)-i|+1$ if $i\leqslant k<j$ (6.21)
and if $j<i$, then let $w(i,j)=w(j,i)$. For any $e\in\mathcal{M}^{s}_{n}(\mathbb{N})$, let
$\mu(e)=\sum_{i=1}^{k}\sum_{j=k+1}^{n}e(i,j)$ (6.22)
and
$\xi(e)=\sum_{i=1}^{n}\sum_{j=i+1}^{n}e(i,j)w(i,j).$ (6.23)
We are ready to prove the following key lemma.
Lemma 6.5.
Suppose that $e\in\mathcal{M}^{s}_{n}(\mathbb{N})$ satisfies property ()(*). Then $\xi(e)\geqslant\mu(e)+k-1$.
Proof.
Let us first consider the case $k=1$. Then $e$ and $w$ are $(2\times 2)$-matrices, in fact
[…]
and
$\xi(e)=\sum_{i=1}^{2}\sum_{j=i+1}^{2}e(i,j)w(i,j)=e(1,2)=\mu(e)=\mu(e)+1-1=\mu(e)+k-1$
as requried. So we can assume that $k>1$.
Denote $\mathcal{S}=\{(i,j)\in\llbracket 1,n\rrbracket^{2}\,:\,i<j\}$ and define $\mathcal{O},\mathcal{X}\subset\mathcal{S}$ to be the sets $\mathcal{O}=\{(i,j)\in\mathcal{S}\,:\,i,j\leqslant k\text{ or }i,j>k\}$, $\mathcal{X}=\mathcal{S}\setminus\mathcal{O}$. We now have
$\xi(e)=\sum_{(i,j)\in\mathcal{S}}e(i,j)w(i,j)\quad\text{ and }\quad\mu(i,j)=\sum_{(i,j)\in\mathcal{X}}e(i,j).$
So it is enough to show that
$\sum_{(i,j)\in\mathcal{S}}e(i,j)w(i,j)\geqslant\sum_{(i,j)\in\mathcal{X}}e(i,j)+k-1.$ (6.24)
Let $\pi\colon\llbracket 1,n\rrbracket\to\llbracket 1,n\rrbracket$ be defined by
[…]
Note that since
$\forall i\in\llbracket 1,n\rrbracket(i\leqslant k\iff f(i)>k)$ (6.25)
$\pi$ is a kind of projection in the sense that it satisfies for all $i\in\llbracket 1,n\rrbracket$
[…]
Let $\rho,\sigma\in\mathcal{M}^{s}_{n}(\mathbb{N})$ be defined by $\rho(i,j)=|\pi(i)-\pi(j)|$ and
$\sigma(i,j)=\begin{cases}0&\text{if }(i,j)\in\mathcal{O}\text{ or }(j,i)\in\mathcal{O}\text{ or }i=j\\ 1&\text{otherwise.}\end{cases}$
Claim 6.5.1.
$w=\rho+\sigma$.
Proof.
We want to show that for all $i,j\in\llbracket 1,n\rrbracket$, $w(i,j)=\rho(i,j)+\sigma(i,j)$. By symmetry it is enough to consider $i\leqslant j$. If $i=j$, then by definitions, $w(i,j)=\rho(i,j)=\sigma(i,j)=0$. So we can assume that $(i,j)\in\mathcal{S}$. Note that
|f(j)i|=|nj+1i|=|ni+1j|=|f(i)j|.|f(j)-i|=|n-j+1-i|=|n-i+1-j|=|f(i)-j|.
The definition of $w$ can be re-written for all $(i,j)\in\mathcal{S}$ as
$w(i,j)=\begin{cases}|j-i|&\text{if }(i,j)\in\mathcal{O}\\ |f(j)-i|&\text{otherwise.}\end{cases}$
Also note that for all $(i,j)\in\mathcal{O}$ we have
|ji|=|nn+11+ji|=|ni+1(nj+1)|=|f(i)f(j)|.|j-i|=|n-n+1-1+j-i|=|n-i+1-(n-j+1)|=|f(i)-f(j)|. (6.27)
We can now easily check the claim case by case:
$(i,j)\in\mathcal{O}\land i,j\leqslant k:$
$\phantom{=}w(i,j)=|j-i|=|\pi(j)-\pi(i)|=\rho(i,j)+0=\rho(i,j)+\sigma(i,j)$
$(i,j)\in\mathcal{O}\land i,j>k:$
[…]
$(i,j)\in\mathcal{X}\land i\leqslant k<j:$
$\phantom{=}w(i,j)=|f(j)-i|+1=|\pi(j)-\pi(i)|+1=\rho(i,j)+1=\rho(i,j)+\sigma(i,j)$
Given two matrices $a$ and $b$, denote by abab their Hadamard product, i.e. the matrix such that $(ab)(i,j)=a(i,j)b(i,j)$. Let us also introduce temporarily the following shorthand notation for a matrix $a$ and a set AA: ${(A)a}=\sum_{(i,j)\in A}a(i,j)$. Then the statement (6.24) becomes
$(\mathcal{S})ew\geqslant(\mathcal{X})e+k-1$ (6.28)
Let $(i_{1},\dots,i_{p})$ be the zigzag $e$-path given by Lemma 6.3. Let $\mathcal{Q}\subset\mathcal{S}$ be defined by $\mathcal{Q}=\mathcal{Q}_{+}\cup\mathcal{Q}_{-}$ where
[…]
By the linearity properties of the sum, Claim 6.5.1, and definitions of $\rho$ and $\sigma$ we have
$(\mathcal{S})ew$ $=(\mathcal{S})e(\rho+\sigma)=(\mathcal{O})e\rho+(\mathcal{X})e(\rho+1)$
$=(\mathcal{O}\setminus\mathcal{Q})e\rho+(\mathcal{O}\cap\mathcal{Q})e\rho+(\mathcal{X}\setminus\mathcal{Q})e(\rho+1)+(\mathcal{X}\cap\mathcal{Q})e(\rho+1)$
$=(\mathcal{O}\setminus\mathcal{Q})e\rho+(\mathcal{O}\cap\mathcal{Q})e\rho+(\mathcal{X}\setminus\mathcal{Q})e\rho+\underline{(\mathcal{X}\setminus\mathcal{Q})e}+(\mathcal{X}\cap\mathcal{Q})e\rho+\underline{(\mathcal{X}\cap\mathcal{Q})e}$
$=(\mathcal{O}\setminus\mathcal{Q})e\rho+(\mathcal{O}\cap\mathcal{Q})e\rho+(\mathcal{X}\setminus\mathcal{Q})e\rho+(\mathcal{X}\cap\mathcal{Q})e\rho+\underline{(\mathcal{X})e}$
$\geqslant(\mathcal{O}\cap\mathcal{Q})e\rho+(\mathcal{X}\cap\mathcal{Q})e\rho+(\mathcal{X})e$
$=(\mathcal{Q}\cap\mathcal{S})e\rho+(\mathcal{X})e.$
At the inequality we just threw away the terms $(\mathcal{O}\setminus\mathcal{Q})e\rho$ and $(\mathcal{X}\setminus\mathcal{Q})e\rho$. The underlined terms are highlighted to make it easier to see which ones were combined to obtain what. In the last step we use the equality $\mathcal{O}\cup\mathcal{X}=\mathcal{S}$ which follows from the definitions of $\mathcal{O}$ and $\mathcal{X}$. So we are left to show that
$(\mathcal{Q}\cap\mathcal{S})e\rho\geqslant k-1.$ (6.29)
We have
$(\mathcal{Q}\cap\mathcal{S})e\rho=(\mathcal{Q}_{+}\cap\mathcal{S})e\rho+(\mathcal{Q}_{-}\cap\mathcal{S})e\rho.$
If $(i,j)\in\mathcal{Q}_{+}\cap\mathcal{S}$, then for some $q\in\llbracket 1,p\rrbracket$, $i=f(i_{q})$, $j=i_{q+1}$, and $f(i_{q})\leqslant i_{q+1}$. In this case
[…]
The inequality follows from Definition 6.2 of a zigzag path. If $(i,j)\in\mathcal{Q}_{-}\cap\mathcal{S}$ then symmetrically for some $q\in\llbracket 1,p\rrbracket$, $j=f(i_{q})$, $i=i_{q+1}$, and $f(i_{q})\geqslant i_{q+1}$. Then by the symmetricity of $e$ and $\rho$,
$e(i,j)\rho(i,j)=e(j,i)\rho(j,i)=e(f(i_{q}),i_{q+1})|\pi(f(i_{q}))-\pi(i_{q+1})|\geqslant|\pi(i_{q})-\pi(i_{q+1})|.$
If $q\in\llbracket 1,p\rrbracket$, then either $(f(i_{q}),i_{q+1})\in\mathcal{Q}_{+}\cap\mathcal{S}$ or $(i_{q+1},f(i_{q}))\in\mathcal{Q}_{-}\cap\mathcal{S}$, but not both, because $\mathcal{S}$ contains exactly one of $(i,j)$ or $(j,i)$ for $i\neq j$. All in all,
$(\mathcal{Q}_{+}\cap\mathcal{S})e\rho+(\mathcal{Q}_{-}\cap\mathcal{S})e\rho\geqslant\sum_{q=1}^{p}|\pi(i_{q})-\pi(i_{q+1})|\geqslant\left|\sum_{q=1}^{p}\pi(i_{q})-\pi(i_{q+1})\right|=|\pi(i_{1})-\pi(i_{p})|.$
But $(i_{1},\dots,i_{p})$ was the zigzag $e$-path given by Lemma 6.3, so $i_{1}=f(1)=n$ and $i_{p}\in\{k,f(k)\}$. By the definition of $\pi$, we have $\pi(i_{1})=1$ and $\pi(i_{p})=k$, so $|\pi(i_{1})-\pi(i_{p})|=|1-k|=k-1$ and we are done. ∎
6.2 Minimal Splitting Curves
We have already defined rays which were used to count certain minimal intersection numbers. In this section we will define a related notion. A splitting curve is an embedding $J\colon I\to\square$ such that $J(t)\in\overset{\circ}{\square}{}\iff t\notin\{0,1\}$. In other words a splitting curve $J$ can be seen as a (complete) tangle diagram $(J,\{1\},\varnothing,\ell_{\varnothing})$ or a knot diagram $(J,\ell_{\varnothing})$ where $\ell_{\varnothing}=\varnothing$ is the empty labeling function. Thus, Definition 4.7 applies also to splitting curves, so we can talk about a splitting curve being in general position with respect to a t-diagram (i.e. they are in mutual general position), or another splitting curve. Also the notation $\operatorname{\mathbf{i}}(\cdot,\cdot)$ of Definition 4.7(4.1) generalizes to splitting curves. Like with t-diagrams, we assume by default that splitting curves are in general position w.r.t. to t-diagrams in each given context.
Definition $6.$
Recall Definition 4.6 of $L_{1}(T),\dots,L_{m}(T)$. Given a t-diagram $T$ and a splitting curve $J$, if the starting point of $J$ is in some $L_{i}(T)$ and the ending point is in some $L_{j}(T)$, then we call $J$ an ijij-splitting curve (of $T$).
Definition 6.7.
Given a t-diagram $T=(T,N,M,\ell)$ denote by $\operatorname{MIN}_{ij}(T)$ the minimum number of intersections that an ijij-splitting curve can have with $T$, formally
$\operatorname{MIN}_{ij}(T)=\min\{\operatorname{\mathbf{i}}(T,J)\,:\,J\text{ is an $ij$-splitting curve of $T$}\}.$ (6.30)
As per our assumption, the curves $J$ in (6.30) range only over those that are in general position w.r.t. $T$. Now define the minimal intersection matrix for a t-diagram $T$, denoted $v_{T}$ to be given by $v_{T}(i,j)=\operatorname{MIN}_{ij}(T)$.
A splitting curve $J$ is called vertical if $J(0)\in H^{+}$ and $J(1)\in H^{-}$, i.e. it passes from the top horizontal edge and goes to the bottom horizontal edge splitting the t-diagram vertically. We introduce a special notation for the vertical minimum intersection number of a t-diagram $T$ defined as
$\operatorname{MIN}_{\operatorname{vert}}(T)=\min\{\operatorname{\mathbf{i}}(T,J)\,:\,J\text{ is a vertical splitting curve of }T\}.$ (6.31)
Again, we note that in this definition we only consider splitting curves in general position with respect to $T$. Recall Definition 4.5 of type of a t-diagram.
Lemma 6.8.
If $T$ is a t-diagram of type $(n_{1},n_{2},n_{3},n_{4})$, then $\operatorname{MIN}_{\operatorname{vert}}(T)\leqslant\min\{n_{2},n_{4}\}$.
Proof.
Suppose w.l.o.g. that $n_{2}\leqslant n_{4}$. Let $\varepsilon$ be so small that the $\varepsilon$-neighborhood $U$ of $V^{-}$ is such that $U\cap T$ is a disjoint union of $n_{2}$ horizontal lines. This is possible without loss of generality by the piecewise linearity assumption. Let $J$ be a vertical straight line inside of $U$. Then $\operatorname{\mathbf{i}}(T,J)=n_{2}$ proving $\operatorname{MIN}_{\operatorname{vert}}(T)\leqslant n_{2}=\min\{n_{2},n_{4}\}$. ∎
Recall that for $(n\times n)$-matrices $a$ and $b$ we denote $a\geqslant b$ if for all $i,j\in\llbracket 1,n\rrbracket$ we have $a(i,j)\geqslant b(i,j)$. We are now ready to state the main lemma of this section which is the key connection between t-diagrams and matrices of Section 6.1.
Lemma 6.9.
Fix $k\in\mathbb{N}_{+}$ and $n=2k$. Suppose that $T=(T,N,M,\ell)$ is a t-diagram of type (1,k1,1,k1)(1,k-1,1,k-1) with $\operatorname{MIN}_{\operatorname{vert}}(T)=k-1$. Let $w$ be the matrix of Definition 6.4. Then $v_{T}\geqslant w$.
The proof is postponed to the end of this section.
Definition 6.10.
Let $T$ and $S$ be t-diagrams, and $\varepsilon>0$. We write $S\subset^{\varepsilon}T$ to mean that there is $C\subset\operatorname{Cross}(T)$ such that the following conditions hold:
  1. (a)
    $|S|\subset|T|\cup\bigcup_{c\in C}\bar{B}(c,\varepsilon)$,
  2. (b)
    For all $c\in C$, $\operatorname{Cross}(S)\cap\bar{B}(c,\varepsilon)=\varnothing$.
We write $S\subset^{\varepsilon}_{\max}T$ if in addition to (a) and (b), all endpoints of $T$ are endpoints of $S$.
Definition 6.11.
A t-diagram $T$ is trivial if $\operatorname{Cross}(T)=\varnothing$. An $\varepsilon$-smoothing of a t-diagram $T$ is a trivial t-diagram $S$ with $S\subset^{\varepsilon}T$. It is maximal if $E(S)=E(T)$, i.e. if $S\subset^{\varepsilon}_{\max}T$. An $m$-bridge for some $m\in\mathbb{N}$ is a trivial t-diagram of type $(0,m,0,m)$ in which every strand connects $V^{-}$ to $V^{+}$. The 0-bridge is the empty t-diagram and is of type $(0,0,0,0)$.
Lemma 6.12.
Suppose $T$ and $S$ are two t-diagrams and $S^{\prime}\subset^{\varepsilon}S$. If $\varepsilon$ is sufficiently small, then $\operatorname{\mathbf{i}}(T,S)\geqslant\operatorname{\mathbf{i}}(T,S^{\prime})$.
Proof.
Suppose $x\in\operatorname{Cross}(S)$. Then by general position there is $\varepsilon_{x}$ such that $B(x,\varepsilon_{x})\cap T=\varnothing$. Let $\varepsilon=\min\{\varepsilon_{x}\,:\,x\in\operatorname{Cross}(S)\}$. The set $\operatorname{Cross}(S)$ is finite so the minimum exists. Suppose $y\in S^{\prime}\cap T$. Since $y\in T$, $y\notin B(x,\varepsilon)$ for all $x\in\operatorname{Cross}(S)$, so by the definition of $\subset^{\varepsilon}$, $y\in S$. Thus $y\in S\cap T$. This shows that $S^{\prime}\cap T\subset S\cap T$ which implies the statement. ∎
Remark 6.13.
It is standard to construct a maximal $\varepsilon$-smoothing of a tangle diagram for sufficiently small $\varepsilon$. This process is illustrated in Figure 5.
Clearly, the relations $\subset^{\varepsilon}$ and $\subset^{\varepsilon}_{\max}$ are transitive for any fixed $\varepsilon$. The relation $S^{\prime}\subset^{\varepsilon}_{\max}S$ does not imply that $S^{\prime}$ is an $\varepsilon$-smoothing of $S$, because an $\varepsilon$-smoothing has no crossings while $S^{\prime}$ can still have them. Nonetheless, there is a transitivity-like relationship between the two notions:
Lemma 6.14.
Suppose $S^{\prime}\subset^{\varepsilon}_{\max}S$ and $S^{*}$ is a maximal $\varepsilon$-smoothing of $S^{\prime}$. Then $S^{*}$ is a maximal $\varepsilon$-smoothing of $S$.
Proof.
It is easy to check from the definitions that $\operatorname{Cross}(S^{\prime})\subset\operatorname{Cross}(S)$. Let $C^{\prime}\subset\operatorname{Cross}(S^{\prime})$ be such that it witnesses $S^{*}\subset^{\varepsilon}S^{\prime}$. Let $C$ witness $S^{\prime}\subset^{\varepsilon}S$. Then $C\cup C^{\prime}$ witnesses $S^{*}\subset^{\varepsilon}S$. By transitivity of $\subset^{\varepsilon}_{\max}$, also $S^{\prime}\subset^{\varepsilon}_{\max}S$. ∎
Lemma 6.15.
If $\beta$ is an $m$-bridge, then $\operatorname{MIN}_{\operatorname{vert}}(\beta)=m$.
Proof.
If $J$ is a vertical splitting curve, then by the Jordan Curve Theorem every strand of $\beta$ must intersect $J$ at least once, so $\operatorname{MIN}_{\operatorname{vert}}(\beta)\geqslant m$. The other direction follows from Lemma 6.8. ∎
Lemma 6.16.
Suppose $T$ is a trivial t-diagram. Then exactly one of the following holds: (1) there is a vertical splitting curve $J$ with $C\cap J=\varnothing$ or (2) there is a strand $s$ of $T$ which connects $V^{-}$ to $V^{+}$.
Proof.
This proof is based on [gale1979hex]. By tilting the square a bit to a rhombus it can be filled with a hexagonal grid so that it constitutes a board for the game of Hex. Assume that the hexagonal grid is so fine that for every hexagon $\sigma$ the set $|T|\cap\operatorname{St}(\sigma)$ is path-connected, where $\operatorname{St}(\sigma)$ is the union of all the hexagons that touch $\sigma$, including $\sigma$ itself. This is possible because $T$ is piecewise linear and therefore locally path connected. Use compactness to find a uniform bound on the size of the hexagons. Note that since $T$ is trivial, $|T|\cap\operatorname{St}(\sigma)$ is in fact a subarc of some strand of $T$. Color the hexagons black if they intersect $T$ and white if they do not. By the Hex Theorem [gale1979hex] there is a white path of hexagons from top to bottom (from $H^{+}$ to $H^{-}$) if and only if there is no black path from left to right (from $V^{-}$ to $V^{+}$). In the former case, we find a vertical splitting curve disjoint from $T$ by following the hexagonal white path from top to bottom. In the latter case let $\sigma_{1},\dots,\sigma_{n}$ be the sequence of black hexagons from $V^{-}$ to $V^{+}$ so that $\sigma_{i}$ touches $\sigma_{i+1}$ for $i\in\llbracket 1,n-1\rrbracket$. For each $i\in\llbracket 1,n\rrbracket$, let $x_{i}\in\sigma_{i}\cap|T|$. By the assumption, for each $i$ there is a subarc of a strand in $T$ connecting $x_{i-1}$, $x_{i}$, and $x_{i+1}$. By the triviality of $T$, all these arcs must be subarcs of the same strand. The union of these subarcs connects $V^{-}$ to $V^{+}$ so it must, in fact, be equal to that strand. ∎
$\Rightarrow Refer to caption$ $\Rightarrow$ Refer to caption $\Rightarrow$ Refer to caption
Figure 5: Illustrating the existence of a maximal $\varepsilon$-smoothing. We start with a t-diagram of type $(1,4,1,4)$ (left). Then all the crossings are replaced by local smoothings. The third diagram from the left is a maximal $\varepsilon$-smoothing (all endpoints are preserved). So is the right-most one which is also loopless (connected components disconnected from the boundary were removed). This figure happens also to illustrate Lemma 6.19. In this case $k=5$, and the circled numbers in the right-most diagram correspond to the clockwise enumeration of $s_{1},\dots,s_{10}$, and the diagram represents $\Sigma_{2}$. For example we can see that $s_{3}$ is connected to $s_{10-3+1}=s_{8}$, as required by Lemma 6.19.
Lemma 6.17.
Suppose $S=(S,N,M,\ell)$ is a t-diagram of type $(0,l_{1},0,l_{2})$ for some $l_{1},l_{2}\in\mathbb{N}$, and $\operatorname{MIN}_{\operatorname{vert}}(S)=m$. Then there is a positive $\varepsilon_{0}$ such that for all positive $\varepsilon<\varepsilon_{0}$ there is an $m$-bridge $\beta$ with $\beta\subset^{\varepsilon}S$. In fact, there is a maximal $\varepsilon$-smoothing $S^{*}$ of $S$ such that $\beta\subset S^{*}$, i.e. every strand of $\beta$ is a strand of $S^{*}$.
Proof.
If $m=0$ then let $\beta=\varnothing$ be the 0-bridge. Clearly $\beta\subset^{\varepsilon}S$ for all $\varepsilon$ and $\beta\subset S^{*}$ for all smoothings $S^{*}$ of $S$, so by Remark 6.13 we are done. Since $T$ is of type $(0,l_{1},0,l_{2})$, the edge $H^{+}$ is outside of the tangle, i.e. $H^{+}\cap S=\varnothing$. Let AA be the connected component of $\square\setminus S$ which contains $H^{+}$, and let $C=\bar{A}\cap S$. For later use, note that it follows that
[…]
We will prove the lemma by induction on $m$ and each step is proved by induction on the cardinality of the set $\operatorname{Cross}(S)\cap C$. So we will prove the following by induction on $\operatorname{Card}(\operatorname{Cross}(S)\cap C)$:
Claim 6.17.1.
Assume that the statement of the lemma holds for $m=p$ and that $\operatorname{MIN}_{\operatorname{vert}}(S)=p+1$. Then there is $\varepsilon_{0}>0$ such that for all positive $\varepsilon<\varepsilon_{0}$ there is an $\varepsilon$-smoothing $S^{*}$ of $S$ and a $(p+1)$-bridge $\beta\subset S^{*}$.
Proof of Claim 6.17.1.
Suppose first that $\operatorname{Cross}(S)\cap C=\varnothing$. Let us show that then $C$ is a union of strands of $S$. For a contradiction assume that there is some strand $s$ of $S$ which is only partially in $C$ and let $x\in C\cap s$ and $y\in s\setminus C$. Take a path $\gamma$ from $H^{+}$ to $x$ which touches $S$ only at $x$. Let $\eta\subset s$ be a path from $x$ to $y$ in $s$ connecting $x$ to $y$. If $\eta$ passes through a crossing, let $z$ be the first crossing $\eta$ encounters on the way from $x$ to $y$. Obtain $\eta^{\prime}$ by shifting $\eta$ slightly in the tubular neighborhood of $s$ in the direction toward $\gamma$, i.e. so that $\eta^{\prime}\cap\gamma$ is non-empty. Obtain $\eta^{\prime\prime}$ by modifying $\eta^{\prime}$ slightly at its other endpoint so that $\eta^{\prime}$ intersects $S$ for the first time at $z$. Then here is a path inside $|\gamma|\cup|\eta^{\prime\prime}|$ which goes from $H^{+}$ to $z$ without intersecting $S$ on the way. By definition of $C$ this means that $z\in C$, a contradiction. Therefore there is no crossing of $S$ on $\eta$. Now shift $\eta$ slightly, again in the direction towards $\gamma$, to obtain $\eta^{\prime}$ so that there is a path in $|\gamma|\cup|\eta^{\prime}|$ which connects $H^{+}$ to $y$, a contradiction again with the assumption that $y\notin C$. Thus, we have that $C$ is in fact a trivial t-diagram consisting of strands of $S$.
We now claim that there is a strand $s$ of $C$ such that $s$ connects $V^{-}$ to $V^{+}$. If not, then by Lemma 6.16 there is a vertical splitting curve $J_{v}$ in the complement of $C$. By the definition of $C$, $J_{v}$ cannot intersect $S$. But then $\operatorname{MIN}_{\operatorname{vert}}(S)=0$, a contradiction (we assumed $\operatorname{MIN}_{\operatorname{vert}}(S)=p+1>p\geqslant 0$).
Let $S_{1}$ be the t-diagram which consists of all other strands of $S$ than $s$. Then $S_{1}\cap s=\varnothing$, for otherwise we would find a crossing in $C$. We claim that $\operatorname{MIN}_{\operatorname{vert}}(S_{1})=p$. Let $J\colon I\to\square$ be a vertical splitting curve with $\operatorname{\mathbf{i}}(S,J)=p+1$. Then at least one of the intersections must be on $s$ by the Jordan Curve Theorem, so $\operatorname{\mathbf{i}}(S_{1},J)\leqslant p$ witnessing $\operatorname{MIN}_{\operatorname{vert}}(S_{1})\leqslant p$. Suppose that $J^{\prime}$ is an arbitrary vertical splitting curve. Let $t\in I$ be the maximum of the set $(J^{\prime})^{-1}[s]$ (the latter set is finite by general position and non-empty by the same argument as above, so it has a maximum). Since $J^{\prime}(t)$ is in $C$, there is a curve $\gamma$ from $H^{+}$ to $J^{\prime}(t)$ which intersects $S$ for the first time at $J^{\prime}(t)$. Let $J^{\prime\prime}=\gamma{}^{\frown}(J^{\prime}\!\restriction\![t,1])$. Then $J^{\prime\prime}$ intersects $s$ exactly once and $S_{1}$ at most as many times as $J^{\prime}$, so $\operatorname{\mathbf{i}}(S_{1},J^{\prime})\geqslant\operatorname{\mathbf{i}}(S_{1},J^{\prime\prime})=\operatorname{\mathbf{i}}(S,J^{\prime\prime})-1\geqslant\operatorname{MIN}_{\operatorname{vert}}(S)-1=p$. By the arbitrary choice of $J^{\prime}$ this means that $\operatorname{MIN}_{\operatorname{vert}}(S_{1})\geqslant p$. We have proved that $\operatorname{MIN}_{\operatorname{vert}}(S_{1})=p$.
By the induction hypothesis there is an $\varepsilon_{0}>0$ such that for all $\varepsilon<\varepsilon_{0}$ there is an $\varepsilon$-smoothing $S^{*}$ of $S_{1}$ and a $p$-bridge $\beta$ such that $\beta\subset S^{*}$. Let $\varepsilon_{0}^{\prime}<\varepsilon_{0}$ be such that an $2\varepsilon^{\prime}_{0}$-neighborhood of $s$ does not contain any crossing of $S_{1}$ (which are the same as crossings of $S$). Let $\varepsilon<\varepsilon^{\prime}_{0}$ and let $\beta$ be a $p$-bridge with $\beta\subset^{\varepsilon}S_{1}$. Let $\beta^{\prime}=\beta\cup s$. Then $\beta^{\prime}$ is a bridge of type $(0,p+1,0,p+1)$, and $\beta^{\prime}\subset^{\varepsilon}S$.
We have proved Claim 6.17.1 in the case where $C$ contains no crossings of $S$. Suppose we have proved this for the case when $C$ contains $q$ crossings of $S$, and suppose that $C$ contains $q+1$ crossings. Fix a vertical splitting curve $J_{0}$ witnessing $\operatorname{MIN}_{\operatorname{vert}}(S)=p+1$. Let $x\in C\cap\operatorname{Cross}(T)$ and let $U$ be a neighborhood of $x$ witnessing the transversality (see Definition 2.1). Select $U$ to be also so small that $U\cap|J_{0}|=\varnothing$. This is possible, because by the assumption of general position, $J_{0}$ does not contain $x$ and $J_{0}$ is compact, so $d(x,J_{0})>0$. Further we can assume that $U\cap\partial\square=\varnothing$. Let $\varepsilon_{0}$ be so small that for all $\varepsilon<\varepsilon_{0}$ we can choose $U=B(x,\varepsilon)$ and also smaller than any $\varepsilon_{0}$ occurred in the previous induction steps. By the assumption that $U$ witnesses transversality, there is a homeomorphism $h\colon U\to[-1,1]^{2}$ such that $h[C]\subset([-1,1]\times\{0\})\cup(\{0\}\times[-1,1])$. Consider $h[A\cap U]$ where AA is as defined in the beginning of the proof of the lemma. By (6.32) there is a point $y\in h[A\cap U]$ such that $y$ is not in any of the coordinate axes. Without loss of generality we may assume that $y$ is in the positive-positive quarter, i.e. $y\in\left]0,1\right[^{2}$. Then we have
$(\{0\}\times[0,1])\cup([0,1]\times\{0\})\subset hC$
Let $S^{\prime}$ be obtained from $S$ by replacing $S\cap U$ with $h^{-1}[\alpha^{+}\cup\alpha^{-}]$ where $\alpha^{+}=[(0,1),(1,0)]$ and $\alpha^{-}=[(-1,0),(0,-1)]$, see Figure 6. Note that one has to reparametrize the immersion of $S^{\prime}$ in an obvious way to make it a t-diagram (see Section 4.5). We did not remove any endpoints in the process, so we have $S^{\prime}\subset^{\varepsilon}_{\max}S$. Let $A^{\prime}$ and $C^{\prime}$ be the sets defined for $S^{\prime}$ in the same way as AA and $C$ were defined for $S$. It remains to show that the induction hypothesis (for $q$) can be applied to $S^{\prime}$ and $C^{\prime}$. Once we do that, for any $\varepsilon<\varepsilon_{0}$ we will obtain a maximal $\varepsilon$-smoothing of $S^{\prime}$ which contains a $(p+1)$-bridge. Since a maximal $\varepsilon$-smoothing of $S^{\prime}$ is a maximal smoothing of $S$ (Lemma 6.14), it will conclude the proof. Thus, we need to show that $\operatorname{Cross}(S^{\prime})\cap C^{\prime}$ has a smaller cardinality than $\operatorname{Cross}(S)\cap C$ and that $\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})=p+1$. First let us prove some preliminary claims.
Claim 6.17.2.
$\left(\,\left]-1,0\right[\times\left]0,1\right[\,\right)\cap hA=\left(\,\left]0,1\right[\times\left]-1,0\right[\,\right)\cap hA=\varnothing$
(Recall that $\left]a,b\right[$ denotes the open interval from $a$ to $b$.)
Proof of Claim 6.17.2.
We will show that $\left]-1,0\right[\times\left]0,1\right[\cap hA=\varnothing$ and the other part follows by symmetry. Suppose for a contradiction that there is $z\in\left]-1,0\right[\times\left]0,1\right[\cap hA$ and let $z^{\prime}=h^{-1}(z)$. Let $y$ be as above, and $y^{\prime}=h^{-1}(y)$. Let $\gamma_{z}$ and $\gamma_{y}$ be paths connecting $z^{\prime}$ and $y^{\prime}$ respectively to $H^{+}$ and let $\varphi$ be the line segment connecting the endpoints of $\gamma_{z}$ and $\gamma_{y}$ in $H^{+}$. Let $\psi$ be the line segment connecting $z$ and $y$. If $\gamma_{z}$ and $\gamma_{y}$ intersect, modify them by perform smoothing operations within AA on their crossings to make them disjoint. Then $h^{-1}[\psi]\cup\gamma_{1}\cup\varphi\cup\gamma_{2}$ is a Jordan curve and it is entirely in $\bar{A}$. Also, the part $\gamma_{1}\cup\varphi\cup\gamma_{2}$ does not intersect $S$. But the remaining part, $h^{-1}[\psi]$ intersects $S$ only in one point. The strand of $S$ which enters the Jordan region must end in $\partial\square$, so this is a contradiction with the Jordan Curve Theorem. ∎
Claim 6.17.3.
$C^{\prime}\setminus U\subset C\setminus U$
Proof of Claim 6.17.3.
Suppose $z\in C^{\prime}\setminus U$ and suppose $\gamma$ is a curve connecting $H^{+}$ to $z$. If $\gamma$ lies outside of $U$, then it witnesses that $z\in C$. Otherwise, w.l.o.g. assume general position between $\gamma$ and $\partial U$, so $\gamma^{-1}[U]$ is a finite union of closed intervals. Let $t_{0},t_{1}]$ be the first one of them (i.e. such that $t_{0}$ is the smallest point with $\gamma(t_{0})\in U$). Then $\gamma\!\restriction\![0,t_{0}]$ witnesses that $\gamma(t_{0})\in A$, so by Claim 6.17.2 $h(\gamma(t_{0}))$ is either in $([0,1]\times\{1\})\cup(\{1\}\times[0,1])$ or in in $([-1,0]\times\{-1\})\cup(\{-1\}\times[-1,0])$. By the Jordan Curve Theorem $t_{1}$ must be in the same of those sets, so by a local ambient isotopy it is possible to remove $\gamma\!\restriction\![t_{0},t_{1}]$ from $U$ thereby reducing the number of components of $\gamma^{-1}[U]$ by one. Inductively we can remove all of them, so we reduce the argument to the one where $\gamma$ is outside of $U$. ∎
By Claim 6.17.3 and the fact that $\operatorname{Cross}(S^{\prime})\setminus U=\operatorname{Cross}(S^{\prime})\setminus\{x\}=\operatorname{Cross}(S)\setminus\{x\}$ we have
$\operatorname{Cross}(S^{\prime})\cap C^{\prime}\setminus U\subset\operatorname{Cross}(S)\cap C\setminus\{x\}$
so $x\in(\operatorname{Cross}(S)\cap C)\setminus(\operatorname{Cross}(S^{\prime})\cap C^{\prime})$. Now let us show that $\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})=p+1$.
Claim 6.17.4.
$A^{\prime}\subset A$.
Proof of Claim 6.17.4.
Use the same argument as in the proof of Claim 6.17.3 to show that $A^{\prime}\setminus U\subset A\setminus U$. Suppose $z\in A^{\prime}\cap U$. Then $h(z)$ cannot be between the lines $\alpha^{+}$ and $\alpha^{-}$ (Figure 6) because then we would be able to find a point in hAhA which contradicts Claim 6.17.2. So $h(z)\in[0,1]^{2}$ or $h(z)\in[-1,0]^{2}$. Suppose $h(z)\in[0,1]^{2}$ and let $\gamma$ be a path connecting $z$ to $H^{+}$. By the same argument as above, $h\left[\gamma\cap U\right]$ cannot pass between $\alpha^{+}$ and $\alpha^{-}$ and also $\gamma\cap S^{\prime}=\varnothing$. But this implies that $\gamma$ in fact witnesses that $z\in A$. The case $h(z)\in[-1,0]^{2}$ follows in the same way. ∎
Recall that $J_{0}$ was defined to witness that $\operatorname{MIN}_{\operatorname{vert}}(S)=p+1$ and lies outside $U$, so also witnesses $\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})\leqslant p+1$. Thus, it is enough to show that $\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})\geqslant p+1$. Suppose $J$ is a vertical splitting curve such that $\operatorname{\mathbf{i}}(S^{\prime},J)=\operatorname{MIN}_{\operatorname{vert}}(S^{\prime})$. If $J$ does not intersect $U$, then it is a splitting curve for $S$, so $\operatorname{\mathbf{i}}(S^{\prime},J)=\operatorname{\mathbf{i}}(S,J)\geqslant p+1$. Otherwise, there is $t$ with $J(t)\in U$. Let $t$ be the maximum such $t$. W.l.o.g. by a general position argument, $J(t)\notin|S^{\prime}|$. If $J(t)\in A^{\prime}$, let $\gamma$ be a curve connecting $H^{+}$ to $J(t)$ which lies in $A^{\prime}$. By Claim 6.17.4 it lies in AA, and so outside $S$. Let $J^{\prime}=\gamma{}^{\frown}(J\!\restriction\![t,1])$.
$\operatorname{\mathbf{i}}(S^{\prime},J)\geqslant\operatorname{\mathbf{i}}(S^{\prime},J^{\prime})=\operatorname{\mathbf{i}}(S,J^{\prime})\geqslant\operatorname{MIN}_{\operatorname{vert}}(S)=p+1.$
Suppose that $J(t)$ is in the middle area, i.e. $h(J(t))$ is between the lines $\alpha^{+}$ and $\alpha^{-}$. Then by Claims 6.17.2 and 6.17.4, $J(t)\notin A^{\prime}$. Thus, there must be at least one $t_{0}<t$ such that $J(t_{0})\in C^{\prime}\subset|S^{\prime}|$, so $\operatorname{\mathbf{i}}(S^{\prime},J\!\restriction\![t_{0},1])+1\leqslant\operatorname{\mathbf{i}}(S^{\prime},J)$ On the other hand there it is not hard to find $\gamma$ which connects $H^{+}$ to $J(t_{0})$ and has exactly one intersection with $S^{\prime}$ and exactly one intersection with $S$. Now
$\operatorname{\mathbf{i}}(S^{\prime},J)\geqslant 1+\operatorname{\mathbf{i}}(S^{\prime}J\!\restriction\![t,1])\geqslant\operatorname{\mathbf{i}}(S^{\prime},\gamma)+\operatorname{\mathbf{i}}(S^{\prime},J\!\restriction\![t,1])=\operatorname{\mathbf{i}}(S^{\prime},\gamma{}^{\frown}(J\!\restriction\![t,1])).$ (6.33)
Since $t$ is the maximum point in $J^{-1}(t)$ and by the choice of $\gamma$, we have
$\operatorname{\mathbf{i}}(S,\gamma{}^{\frown}(J\!\restriction\![T,1]))\geqslant\operatorname{MIN}_{\operatorname{vert}}(S)=p+1.$
Suppose now that $t$ is such that $h(J(t))$ is below $\alpha^{-}$ and that area is not in $A^{\prime}$. But then the component of $J^{-1}[U]$ to which $t$ belongs must be either contained entirely to that area, in which case we can remove it by a local ambient isotopy out from $U$ and reduce the number of components in $J^{-1}[U]$ (similarly as we did in the proof of Claim 6.17.3), or it intersects $\alpha^{-}$ in which case we apply a similar argument as in (6.33) replacing “1+1+” with “2+2+”. ∎
This completes the proof of the lemma. ∎
Figure 6: Illustration of the local smoothing performed in the proof of Claim 6.17.1 of Lemma 6.17.
Lemma 6.18.
Suppose $T$ is a t-diagram of type $(0,k,0,k)$ and $\operatorname{MIN}_{\operatorname{vert}}(T)\geqslant k-1$. Then $\operatorname{MIN}_{\operatorname{vert}}(T)\geqslant k$.
Proof.
Let $J$ be a vertical splitting curve which has the minimum intersection number with $T$. Assume for a contradiction that $\operatorname{\mathbf{i}}(T,J)=k-1$. Let $\varepsilon$ be so small that
$J\cap\bigcup_{x\in\operatorname{Cross}(T)}B(x,\varepsilon)=\varnothing$
and so that the conclusion of Lemma 6.17 holds for $T$. Then there is a maximal $\varepsilon$-smoothing $T^{*}$ of $T$ and a (k1)(k-1)-bridge $\beta\subset T^{*}$ which has k1k-1 strands. But there is one more endpoint on $V^{-}$ and one on $V^{+}$ which are not connected by strands of $\beta$. So there must be a strand $s$ in $T^{*}$ connecting them. Let $\beta^{\prime}=\beta\cup s$. Then $\beta^{\prime}$ is a $k$-bridge. Since $J$ goes outside of a neighborhood where $T$ has been modified, $\operatorname{\mathbf{i}}(T,J)=\operatorname{\mathbf{i}}(T^{*},J)\geqslant\operatorname{\mathbf{i}}(\beta^{\prime},J)\geqslant\operatorname{MIN}_{\operatorname{vert}}(\beta^{\prime})=k$ where the last equality is by Lemma 6.15. A contradiction. ∎
Recall the function f(j)=nj+1f(j)=n-j+1 from Section 6.1.
Lemma 6.19.
Let $k\in\mathbb{N}_{+}$ and $n=2k$. Let $T$ be a t-diagram of type (1,k1,1,k1)(1,k-1,1,k-1) such that $\operatorname{MIN}_{\operatorname{vert}}(T)=k-1$. Let $t_{1},\dots,t_{n}$ be the endpoints of $T$ enumerated counterclockwise starting from the only point on $H^{+}$. Let $s_{1},\dots,s_{n}$ be the same points enumerated clockwise starting from the same point, i.e. $s_{1}=t_{1}$ and $s_{i}=t_{n-i+2}$ for all $i\in\llbracket 2,n\rrbracket$. Then there is $\varepsilon_{0}$ such that for all $\varepsilon<\varepsilon_{0}$ there are two trivial t-diagrams $\Sigma_{1}$ and $\Sigma_{2}$ satisfying the following conditions.
  1. 1.
    The $i$-th strand of $\Sigma_{1}$ connects $t_{i}$ to $t_{f(i)}$ for all $i\in\llbracket 1,k\rrbracket$
  2. 2.
    The $i$-th strand of $\Sigma_{2}$ connects $s_{i}$ to $s_{f(i)}$ for all $i\in\llbracket 1,k\rrbracket$
  3. 3.
    $\Sigma_{1},\Sigma_{2}\subset^{\varepsilon}_{\max}T$.
Remark 6.20.
Figures 5 and 7 illustrate this lemma.
Proof.
Clearly proving the existence of $\Sigma_{1}$ implies the existence of $\Sigma_{2}$ by mirroring $T$ along the middle vertical axis. Note that by the definition of type (1,k1,1,k1)(1,k-1,1,k-1) and the counterclockwise enumeration of $t_{i}$, $t_{1}$ is the only point on $H^{+}$ and $t_{k+1}$ is the only point on $H^{-}$. There is an orientation-preserving self-homeomorphism $h$ of $\square$ taking $t_{1}$ to $V^{-}$, $t_{k+1}$ to $V^{+}$, fixing all other points $t_{i}$ as well as the top-right corner and lower-left corners of $\square$. In particular
$h^{-1}H^{\pm}\subset H^{\pm}.$ (6.34)
For any vertical splitting curve $J$ we have $\operatorname{\mathbf{i}}(hT,J)=\operatorname{\mathbf{i}}(T,h^{-1}J).$ If $J$ is a vertical splitting curve, then so is $h^{-1}J$, by (6.34). Thus, $\operatorname{MIN}_{\operatorname{vert}}(hT)\geqslant\operatorname{MIN}_{\operatorname{vert}}(T)=k-1$ and hThT is of type $(0,k,0,k)$. By Lemmas 6.18 and 6.8 we have $\operatorname{MIN}_{\operatorname{vert}}(hT)=k$. Let $\beta$ be the $k$-bridge given by Lemma 6.17. Then $\Sigma_{1}=h^{-1}\beta$ is as desired. ∎
For a sequence $a_{1},\dots,a_{l}\in\partial\square$ with $l\geqslant 3$ denote by $\mathop{\bm{\circlearrowleft}}\!\left(a_{1},\dots,a_{l}\right)$ the statement that these points are arranged counterclockwise.
Fact 6.21.
For splitting curves $J$ and $\sigma$ with $\mathop{\bm{\circlearrowleft}}\!\left(J(0),\sigma(0),J(1),\sigma(1)\right)$ we have $\operatorname{\mathbf{i}}(\sigma,J)\geqslant 1$.
Proof.
Direct application of [Bi83, Theorem III.3.A]. ∎
We are now ready to prove Lemma 6.9.
Proof of Lemma 6.9.
Let $J$ be an ijij-splitting curve for $T$. W.l.o.g. $i\leqslant j$. If $i=j$, $w(i,j)=0$, so the statement is trivial. So we are left with the case $i<j$. Suppose first that $i<j\leqslant k$. Since $T$ is of type (1,k1,1,k1)(1,k-1,1,k-1), $L_{i}(T)$ and $L_{j}(T)$ are now subsets of $V^{-}$. We have to show that $\operatorname{\mathbf{i}}(T,J)\geqslant w(i,j)=|j-i|=j-i$. Suppose not. Let $J_{1}$ be the line from the top endpoint of $V^{-}$, $(0,1)$, to $J(0)$. Then $\operatorname{\mathbf{i}}(T,J_{1})=i-1$. Let $J_{2}$ be the line from $J(1)$ to the lower endpoint of $V^{-}$, $(0,0)$. Then $\operatorname{\mathbf{i}}(T,J_{2})=k-j$. Let $J_{*}=J_{1}{}^{\frown}J{}^{\frown}J_{2}$. Then $\operatorname{\mathbf{i}}(T,J_{*})=\operatorname{\mathbf{i}}(T,J_{1})+\operatorname{\mathbf{i}}(T,J)+\operatorname{\mathbf{i}}(T,J_{2})=i-1+\operatorname{\mathbf{i}}(T,J)+k-j<i-1+j-i+k-j=k-1.$ By pushing $J_{*}$ slightly away from $V^{-}$ into $\square$, obtain a vertical splitting curve $J_{*}^{\prime}$ with $\operatorname{\mathbf{i}}(T,J_{*}^{\prime})<k-1$, a contradiction. Same argument for $i>j>k$ noting that then $L_{i}(T),L_{j}(T)\subset V^{+}$. Suppose
$i\leqslant k<j.$ (6.35)
In this case $i\in V^{-}$ and $j\in V^{+}$. We consider separately the cases $f(j)\leqslant i$ and $f(j)>i$. Suppose first that
$f(j)\leqslant i$ (6.36)
Figure 7 illustrates the forthcoming technicalities. Let $\Sigma_{1}\subset^{\varepsilon}_{\max}T$ be as given by Lemma 6.19. Let $\sigma_{1},\dots,\sigma_{k}$ be the strands of $\Sigma_{1}$ such that $\sigma_{l}$ connects $t_{l}$ to $t_{f(l)}$, i.e., we can think of $\sigma_{l}$ as a splitting curve $\sigma_{l}\colon I\to\square$ with
$\sigma_{l}(0)=t_{l}\quad\text{ and }\quad\sigma_{l}(1)=t_{f(l)}$ (6.37)
where $t_{1},\dots,t_{n}$ are also as given in the formulation of Lemma 6.19. Note that
$\mathop{\bm{\circlearrowleft}}\!\left(t_{1},\dots,t_{n}\right).$ (6.38)
By assuming that $\varepsilon$ is sufficiently small and using the fact that $\Sigma_{1}\subset^{\varepsilon}T$, by Lemma 6.12 we can assume that $\operatorname{\mathbf{i}}(T,J)\geqslant\operatorname{\mathbf{i}}(\Sigma_{1},J)$, so it is enough to bound the latter. Let
$A=\{l\in\llbracket 1,k\rrbracket\,:\,\mathop{\bm{\circlearrowleft}}\!\left(\sigma_{l}(0),J(0),\sigma_{l}(1),J(1)\right)\}$
Then by Fact 6.21
$\operatorname{\mathbf{i}}(\Sigma_{1},J)\geqslant\sum_{l\in A}\operatorname{\mathbf{i}}(\sigma_{l},J)\geqslant\operatorname{Card}(A).$
So it is enough to compute a lower bound for $\operatorname{Card}(A)$. By (6.37),
$A=\{l\in\llbracket 1,k\rrbracket\,:\,\mathop{\bm{\circlearrowleft}}\!\left(t_{l},J(0),t_{f(l)},J(1)\right)\}$
By definition of $L_{i}(T)$ and $L_{j}(T)$, we have that $\mathop{\bm{\circlearrowleft}}\!\left(t_{i},J(0),t_{i+1}\right)$ and $\mathop{\bm{\circlearrowleft}}\!\left(t_{j},J(1),t_{j+1}\right)$, so
$\mathop{\bm{\circlearrowleft}}\!\left(t_{l},J(0),t_{f(l)},J(1)\right)$ (6.39)
holds if $l\leqslant i$ and $f(l)\leqslant j$. The latter is equivalent to $f(j)\leqslant l$ by (6.2). Thus, (6.39) holds if $l\in\llbracket 1,i\rrbracket\cap\llbracket f(j),k\rrbracket=\llbracket f(j),i\rrbracket$, thus
[…]
For the case $f(j)>i$ use $\Sigma_{2}$ and $s_{1},\dots,s_{n}$ in a symmetric fashion. ∎
Figure 7: Illustration of Lemma 6.19 and the proof of Lemma 6.9. The gray lines are strands of $\Sigma_{1}$. In this example $k=3$ and $n=6$. It is evident from the picture that if, say $i=3$ and $j=5$, then any splitting curve going from $L_{i}(T)$ to $L_{j}(T)$ must intersect at least |f(j)i|+1=|65+13|+1=2|f(j)-i|+1=|6-5+1-3|+1=2 strands of $\Sigma_{1}$.
6.3 Bounding the Number of Crossings Between Overlapping Tangle Diagrams
Fix $k\in\mathbb{N}_{+}$ and $n=2k$. Next we will define the matrix $e(T,S)$ for particular t-diagrams $T$ and $S$. Viewed as a multigraph, the vertices of $e(T,S)$ are the sets $L_{1}(T),\dots,L_{n}(T)$ and the edges are the strands of $S^{*}$ where $S^{*}$ is as given by Lemma 6.17 for $S$.
Definition 6.22.
Suppose that $T$ is a t-diagram of type (1,k1,1,k1)(1,k-1,1,k-1) and $S$ a t-diagram of type $(0,l_{1},0,l_{2})$ for some $l_{1},l_{2}\in\mathbb{N}$. We will define the $(n\times n)$-matrix $e(T,S)$ as follows. Let $\varepsilon_{0}$ be as given by Lemma 6.17 for $S$ and let $\varepsilon<\varepsilon_{0}$ so small that for any $x,y\in\operatorname{Cross}(T\times S)$, $B(x,\varepsilon)$ and $B(y,\varepsilon)$ are disjoint and witness the transversality of the respective crossing. Let $S^{*}$ be an $\varepsilon$-smoothing $S$ such that $S^{*}$ is a trivial t-diagram and there is a trivial bridge of type $(0,m,0,m)$ as a subset of $S^{*}$ where $m=\operatorname{MIN}_{\operatorname{vert}}(S)$. Such smoothing exists by Lemma 6.17 and the choice of $\varepsilon$ above. For any $(i,j)\in\llbracket 1,n\rrbracket^{2}$, if $i\neq j$, define $e(T,S)(i,j)$ to be the number of strands of $S^{*}$ that connect $L_{i}(T)$ to $L_{j}(T)$. Define $e(T,S)(i,i)$ to be $2x$ where $x$ is the number of strands which start and end at $L_{i}(T)$.
Remark 6.23.
A careful observation will reveal that $e(T,S)$ is not unique, because it depends on the choice of $S^{*}$. But this does not matter to us. Any choice of $e(T,S)$ will work for our purposes as long as it is obtained from any such $S^{*}$.
Proposition 6.1 (Overlapping t-Diagrams).
Let $l_{1},l_{2}\in\mathbb{N}$. Suppose that $T$ and $S$ are t-diagrams of types (1,k1,1,k1)(1,k-1,1,k-1) and $(0,l_{1},0,l_{2})$ respectively. Suppose that $e(T,S)$ satisfies property ()(*) (Definition 6.1). Then $\operatorname{\mathbf{i}}(T,S)\geqslant\operatorname{MIN}_{\operatorname{vert}}(S)+k-1.$
Proof.
Let $\varepsilon$ and $S^{*}$ be as in Definition 6.22. In particular, $S^{*}$ contains an $m$-bridge for $m=\operatorname{MIN}_{\operatorname{vert}}(S)$. By Lemma 6.12, $\operatorname{\mathbf{i}}(T,S)\geqslant\operatorname{\mathbf{i}}(T,S^{*})$. Since $S^{*}$ and $S$ have the same number of endpoints, they have the same number of strands which is equal to $l=(l_{1}+l_{2})/2$. Let $\mathscr{S}$ be the set of strands of $S^{*}$, $\mathscr{S}=\{s_{1},\dots,s_{l}\}$. Then
$\operatorname{\mathbf{i}}(T,S^{*})\geqslant\sum_{s\in\mathscr{S}}\operatorname{\mathbf{i}}(T,s).$
(The inequality may be strict, if there are loops in $S^{*}$ which are not counted in the sum.)
For all $(i,j)\in\llbracket 1,n\rrbracket^{2}$ let $\mathscr{S}_{ij}\subset\mathscr{S}$ be the set of those strands which go from $L_{i}(T)$ to $L_{j}(T)$. Then $\{\mathscr{S}_{ij}\}_{1\leqslant i<j\leqslant n}$ is a partition of a (possibly proper) subset of $\mathscr{S}$ (we are not counting strands in $\mathscr{S}_{ii}$). Therefore
$\operatorname{\mathbf{i}}(T,S^{*})\geqslant\sum_{i=1}^{n}\sum_{j=i+1}^{n}\sum_{s\in\mathscr{S}_{ij}}\operatorname{\mathbf{i}}(T,s).$
For each $s\in\mathscr{S}_{ij}$ we have $\operatorname{\mathbf{i}}(T,s)\geqslant v_{T}(i,j)$ by Definition 6.7 of $v_{T}$. On the other hand, $\operatorname{Card}(\mathscr{S}_{ij})=e(T,S)(i,j)$ by the definition of $e(T,S)$, so by Lemma 6.9,
$\operatorname{\mathbf{i}}(T,S^{*})\geqslant\sum_{i=1}^{n}\sum_{j=i+1}^{n}e(T,S)v_{T}(i,j)\geqslant\sum_{i=1}^{n}\sum_{j=i+1}^{n}e(T,S)w(i,j)=\xi(e(T,S)).$
By the assumption that $e(T,S)$ satisfies property ()(*) and by Lemma 6.5 we now have
$\operatorname{\mathbf{i}}(T,S^{*})\geqslant\mu(e(T,S))+k-1.$ (6.40)
So it remains to show that $\mu(e(T,S))\geqslant m=\operatorname{MIN}_{\operatorname{vert}}(S)$. By definition,
$\mu(e(T,S))=\sum_{i=1}^{k}\sum_{j=k+1}^{n}e(T,S)(i,j)=\sum_{i=1}^{k}\sum_{j=k+1}^{n}\operatorname{Card}(\mathscr{S}_{ij}).$
The latter quantity counts exactly the number of strands of $S^{*}$ which go from $V^{-}$ to $V^{+}$. Since $S^{*}$ contains an $m$-bridge, it is at least $m$. ∎
Figure 8: An $R^{*}_{0}$-move keeping a neighborhoods of $o$, $o^{-}$, and $o^{+}$ fixed.
7 Bounding the Number of Hybrid Crossings
As mentioned in the beginning of Section 6, in this section we will give a lower bound for $\operatorname{\mathbf{i}}(\widehat{K}^{-},\widehat{K}^{+})$ which will be the missing piece for the proof of Theorem 3.1.
7.1 Subdivided Tangle Diagrams
In order to state the Fan Lemma 7.7 in a sufficiently general form so that it can later be used for links and not only for knots, we need to extend the definitions of the partial knot diagrams $K^{\pm}$ to tangle diagrams more generally.
Definition 7.1 (Subdivided Tangle Diagram).
A tuple $(T,T^{+},T^{-},N,M,\ell)$ is a subdivided t-diagram if the following hold:
  1. ($PT_{1}$)
    $(T,N,M,\ell)$ is a t-diagram of type $(1,0,1,0)$
  2. ($PT_{2}$)
    $N=\{1\}$ is a singleton
  3. ($PT_{3}$)
    There is a partition $M=M_{+}\cup M_{-}$ such that $T^{+}=T\!\restriction\!\big((I^{+}\times\{1\})\cup(S^{1}\times M_{+})\big)$ and $T^{-}=T\!\restriction\!\big((I^{-}\times\{1\})\cup(S^{1}\times M_{-})\big)$.
A subdivided t-diagram $(T,T^{+},T^{-},\{1\},M,\ell)$ is centered if the knot diagram defined by $T\!\restriction\!(I\times\{1\})$ is centered. The lower and upper minimal splitting numbers $\operatorname{MIN}_{\pm}(T^{\mp})$ are defined in the same way as for (partial) knot diagrams. Consistent with the terminology for knot diagrams, call $T^{\pm}$ the partial tangle diagrams, or partial t-diagrams. The $\sim^{*}_{\mathcal{R}}$, $\sim^{*}_{0}$, and $\sim_{\mathcal{R}}$-equivalences are extended to subdivided t-diagrams in an obvious way. A subdivided t-diagram is permitted if for all $x\in T^{+}\cap T^{-}$, $\ell^{-1}(x)\in(I^{+}\cup(S^{1}\times M_{+}))$. Extend the definitions of hThT and $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$ (see Definition 4.4 and Section 4.4) for partial and subdivided tangle diagrams in an obvious way, but note that we will treat $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$ as a t-diagram, not as a subdivided one, because it might not be of type $(1,0,1,0)$ as condition ($PT_{1}$) above requires. Also, if $T^{+}$ is a partial t-diagram, then $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$ is a t-diagram when $o\notin G$. All situations where we will use this notion will be “non-problematic” in this sense.
7.2 Upper and Lower Splitting Curves
Definition 7.2.
A lower splitting curve is an embedding $J^{-}\colon\left[0,\frac{1}{2}\right]\to\square^{-}$ such that $J^{-}(0)=o^{-}$ and $J^{-}\left(\frac{1}{2}\right)=o$. An upper splitting curve is symmetrically an embedding $J^{+}\colon\left[\frac{1}{2},1\right]\to\square^{+}$ such that $J^{+}\left(\frac{1}{2}\right)=o$ and $J^{+}(1)=o^{+}$. Given a t-diagram $T$, define the lower splitting number of $T^{+}$ by
$\operatorname{MIN}_{-}(T^{+})=\min\{\operatorname{\mathbf{i}}(T^{+},J^{-})-1\,:\,J^{-}\text{ is a lower splitting curve}\}$
and symmetrically, define the upper splitting number of $T^{-}$ by
$\operatorname{MIN}_{+}(T^{-})=\min\{\operatorname{\mathbf{i}}(T^{-},J^{+})\,:\,J^{+}\text{ is an upper splitting curve}\}.$
As in the definition of $\operatorname{MIN}_{\operatorname{ray}}$, note that the point $o$ is not counted in $\operatorname{\mathbf{i}}(T^{\pm},J^{\mp})$
Lemma 7.3.
For a subdivided t-diagram $T$, for all $\pm,\mp\in\{+,-\}$, $\operatorname{MIN}_{\pm}(T^{\mp})\geqslant\operatorname{MIN}_{\operatorname{ray}}(T^{\mp})$.
Proof.
All upper and lower splitting curves are rays. ∎
7.3 Straightening Rays
Let $v^{-}=\left(0,\tfrac{1}{2}\right)$ and $v^{+}=\left(1,\tfrac{1}{2}\right)$ be midpoints of $V^{-}$ and $V^{+}$ respectively. Denote by $H_{\tfrac{1}{2}}=[0,1]\times\left\{\tfrac{1}{2}\right\}=[v^{-},v^{+}]$ the middle horizontal line segment. Let $b_{\pm}(\delta)$ be the point $o\pm(\delta/2,0)$. For $r>0$, let $\square(r)$ be the square with side-length $r$ and center $o$:
$\square(r)=\left[\frac{1-r}{2},\frac{1+r}{2}\right]^{2}.$
For instance, $\square=\square(1)$.
$\Rightarrow$
Figure 9: Illustration of Lemma 7.4. The white area in (a) is bounded by the Jordan curve $C^{+}_{1}\cup J_{1}\cup C^{+}_{\delta}\cup J_{2}$. It is mapped by a canonical homeomorphism to the Jordan curve $C^{+}_{1}\cup[v^{-},b_{-}(\delta)]\cup C^{+}_{\delta}\cup[b_{+}(\delta),v^{+}]$ in (b). This homeomorphism extends to the white area by the Jordan Curve Theorem. The light gray area in (a) is bounded by $C^{-}_{1}\cup J_{1}\cup C^{-}_{\delta}\cup J_{2}$ and is mapped to the corresponding area in (b). Together these homeomorphisms extend to a self-homeomorphism of the square which fixes the light red area (which is $\square(\delta)$ pointwise. One way to visualize this homeomorphism is to imagine an isotopy which rotates the light red square $360^{\circ}$ clockwise.
Lemma 7.4.
Suppose $0<\delta<\frac{1}{8}$. Suppose $J_{1}$, $J_{2}\colon I\to\square(1-\delta)$ are embeddings such that $J_{1}\cap J_{2}=\varnothing$, $J_{1}(0)=b_{-}(\delta)$, $J_{1}(1)=v^{-}$, $J_{2}(0)=b_{+}(\delta)$, $J_{2}(1)=v^{+}$, and for $t\in\overset{\circ}{I}{}$, $J_{l}(t)\in\overset{\circ}{\square}{}\setminus\square(\delta)$, $l\in\{1,2\}$. Then there is a homeomorphism $h\colon\square\to\square$ such that $|h\circ J_{1}|=[v^{-},b_{-}(\delta)]$, $|h\circ J_{2}|=[b_{+}(\delta),v^{+}]$ and $h$ is identity on $\partial\square\cup\square(\delta)$.
Proof.
For $r\leqslant 1$, let $C^{\pm}_{r}$ be the upper and lower halves of $\partial\square(r)$, $C^{\pm}_{r}=\partial\square(r)\cap\square^{\pm}$. Then $C^{\pm}_{\delta}$ are curves connecting $b_{-}(\delta)$ to $b_{+}(\delta)$ and $C^{\pm}_{1}$ are curves connecting $v^{-}$ to $v^{+}$. Now $C^{+}_{1}\cup J_{1}\cup C^{+}_{\delta}\cup J_{2}$ and $C^{-}_{1}\cup J_{1}\cup C^{-}_{\delta}\cup J_{2}$ are Jordan curves. Let $f_{1}\colon|J_{1}|\to[v^{-},b_{-}(\delta)]$ be given by $f_{1}(J_{1}(t))=(1-t)v^{-}+tb_{-}(\delta)$. Similarly let $f_{2}\colon|J_{2}|\to[b_{+}(\delta),v^{+}]$ be given by $f_{2}(J_{2}(t))=(1-t)b_{+}(\delta)+tv^{+}$. Denote by $\Gamma^{\pm}_{r}$ the identity map on $C^{\pm}_{r}$. Then $\Gamma^{\pm}_{1}\cup f_{1}\cup f_{2}\cup\Gamma^{\pm}_{\delta}$ is a homeomorphism between two Jordan curves, so it extends to a homeomorphism $h^{\pm}$ between the bounded regions. Let $h_{0}=h^{+}\cup h^{-}$. This homeomorphism is identity on $C^{+}_{\delta}\cup C^{-}_{\delta}=\partial\square(\delta)$ and on $C_{1}^{+}\cup C^{-}_{1}=\partial\square$. So it extends to a self-homeomorphism $h$ of $\square$ which keeps $\partial\square\cup\square(\delta)$ fixed pointwise. See Figure 9 for an illustration. ∎
Denote $d_{\pm}(\delta)=o\pm((1-\delta)/2,0)=b_{\pm}(1-\delta)$. The following follows by simple rescaling and the assumption $\delta<1/8$:
Corollary 6.
Lemma 7.4 applies with $\square$ replaced by $\square(1-\delta)$ and $v^{\pm}$ replaced by $d_{\pm}(\delta)$. ∎
For $0<r<1$, let
$U(r)=\overset{\circ}{\square}{}\setminus\square(1-r).$
Clearly $U(r)$ is open and its boundary is $\partial\square(1-r)\cup\partial\square$.
Lemma 7.5 (Straightening).
Suppose that $J\colon I\to\square$ is a ray such that $J(1)=v^{+}$ and for some $0<\delta<\frac{1}{8}$,
$J\cap\square(2\delta)=[0,b_{+}(2\delta)]\quad\text{ and }\quad J\cap U(2\delta)=[d_{+}(2\delta),v^{+}].$
Then there is a boundary-fixing self-homeomorphism $h$ of $\square$ such that $|h\circ J|=[o,v^{+}]$ and $h$ is the identity on $U(\delta)\cup\square(\delta)$.
Proof.
We will work in $\square(1-\delta)$ and show that there is a self-homeomorphism $h$ of $\square(1-\delta)$ which fixes $\square(\delta)\cup\partial\square(1-\delta)$ and straightens $J$. The lemma then follows by extending $h$ by identity to $\square\setminus\square(1-\delta)$. W.l.o.g. assume that $J(1-\delta)=d_{+}(\delta)$. Using a tubular neighborhood of $J$ find a parallel curve $J^{\prime}$ such that $J^{\prime}(0)\in\partial\square(2\delta)$, $J^{\prime}(1)\in\partial\square(1-2\delta)$, and $J^{\prime}\cap J=\varnothing$. Then $J^{\prime}(1)\in\partial U(2\delta)$.
The set $A=\square(2\delta)\setminus(\square(\delta)\cup J)$ is connected and $b_{-}(\delta)$ is on the boundary of AA. So there is $\gamma_{1}\subset A\cup\{b_{-}(\delta)\}$ which starts at $b_{-}(\delta)$ and ends at $J^{\prime}(0)$. The set $B=U(2\delta)\setminus(U(\delta)\cup J)$ is also connected and $d_{-}(\delta)$ is on the boundary of $B$, so there is a curve $\gamma_{2}\subset B\cup\{d_{-}(\delta\}$ which starts at $J^{\prime}(1)$ and ends at $d_{-}(\delta)$. Let $J_{1}=J\!\restriction\![0,1-\delta]$ and let
$J_{2}=[o,b_{-}(\delta)]\cup\gamma_{1}\cup J^{\prime}\cup\gamma_{2}$
Then $J_{1}$ and $J_{2}$ satisfy the assumptions of Lemma 7.4 for $\square(1-\delta)$ instead of $\square$. So by Corollary 6 there is a required homeomorphism. ∎
Lemma 7.6.
Let $T$ be a centered subdivided t-diagram. Let $J_{0}$ be the ray $J_{0}(t)=(1-t)o+tv^{+}$. Then there is an $\widehat{T}\sim^{*}_{0}T$ such that
$\operatorname{\mathbf{i}}(J,\widehat{T}^{-})=\operatorname{MIN}_{\operatorname{ray}}(T^{-})=\operatorname{MIN}_{\operatorname{ray}}(\widehat{T}^{-}).$
Proof.
Let $J$ be a ray such that $\operatorname{\mathbf{i}}(J,T^{-})=\operatorname{MIN}_{\operatorname{ray}}(T^{-})$. Let $\delta>0$ be such that $B(o,\delta)\cap T^{-}$ is one line segment and also $U(\delta)\cap T^{-}$ is one line segment. Just in case assume also that $\delta<1/8$ so that $\square_{\delta}$ and $U(\delta)$ are disjoint. Let $t$ be the maximum of the set $J^{-1}[\square_{\delta}]$. Let $\gamma$ be a curve connecting $b_{+}(\delta)$ to $J(t)$ inside $(\square_{\delta}\setminus T^{-})\cup\{b_{+}(\delta)\}$. Let $J^{\prime}=[o,b_{+}(\delta)]\cup\gamma\cup J\!\restriction\![\delta,1]$. Then $J^{\prime}$ is still a ray witnessing $\operatorname{MIN}_{\operatorname{ray}}(T^{-})$. Let $t^{\prime}$ be the minimum of $(J^{\prime})^{-1}[U(\delta)]$. Let $\gamma^{\prime}$ be a curve in $(U(\delta)\setminus T^{-})\cup\{v^{+}\}$ connecting $J^{\prime}(t^{\prime})$ to $v^{+}$. Let $J^{\prime\prime}=(J^{\prime}\!\restriction\![0,t^{\prime}])\cup\gamma^{\prime}$. Then $J^{\prime\prime}$ still witnesses $\operatorname{MIN}_{\operatorname{ray}}(T^{-})$. After reparametrization one can make sure that $J^{\prime\prime}$ satisfies the assumptions of the Straightening Lemma 7.5, so let $h$ be the homeomorphism given by that Lemma. Then $h$ witnesses an $R^{*}_{0}$-move which results in $\widehat{T}=hT$ and $hJ=[0,v^{+}]$ witnesses $\operatorname{MIN}_{\operatorname{ray}}(T^{-})=\operatorname{MIN}_{\operatorname{ray}}(\widehat{T}^{-})$ by Lemma 5.2. ∎
Figure 10:
7.4 The Fan Lemma
Recall the notions of axis-aligned rectangle and homeomorphism, a $G$-tangle, $T\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G$ from Definitions 4.3 and 4.4 and their adaptations to subdivided t-diagrams from Definition 7.1. Fix $G$ to be the square
$G=\left[\tfrac{3}{8},\tfrac{5}{8}\right]\times\left[\tfrac{1}{8},\tfrac{3}{8}\right]$
Note that it is axis-aligned (Definition 4.1). Fix also $g$ to be the unique axis-aligned homeomorphism $g\colon G\to\square$ (cf. Definition 4.2). Let us introduce the notation $\chi_{h}(T)=\operatorname{\mathbf{i}}(T^{-},T^{+})$ to denote the number of “hybrid” crossings in a tangle diagram.
Lemma 7.7 (The Fan Lemma).
Suppose that $T=(T,T^{+},T^{-},\{1\},M,\ell)$ is a centered subdivided t-diagram with one strand and $M$ loops. Then there is a centered subdivided t-diagram $\operatorname{Fan}(T)\sim^{*}_{0}T$,
$\operatorname{Fan}(T)=(\operatorname{Fan}(T),\operatorname{Fan}(T)^{+},\operatorname{Fan}(T)^{-},\{1\},M,\operatorname{Fan}(\ell)),$
with the following properties. The diagrams $\operatorname{Fan}(T)$, $\operatorname{Fan}(T)^{+}$, and $\operatorname{Fan}(T)^{-}$ are in general position with respect to $G$. Denote $S^{-}=g(\operatorname{Fan}(T)^{-}\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G)$ and $S^{+}=g(\operatorname{Fan}(T)^{+}\mathbin{\vtop{\halign{#\cr\hfil$\cap$\hfil\cr\hfil$\cdot$\hfil\cr}}}G)$. Then:
  1. (F1F{1})
    $\operatorname{Cross}(\operatorname{Fan}(T))\subset G$ and $\operatorname{Cross}_{h}(\operatorname{Fan}(T))\subset G$
  2. (F2F{2})
    $\operatorname{MIN}_{+}(\operatorname{Fan}(T)^{-})=\operatorname{MIN}_{\operatorname{ray}}(T^{-})=\operatorname{MIN}_{\operatorname{ray}}(\operatorname{Fan}(T)^{-})$
  3. (F3F{3})
    $\operatorname{MIN}_{\operatorname{vert}}(S^{-})=\operatorname{MIN}_{+}(\operatorname{Fan}(T)^{-})$
  4. (F4F{4})
    $S^{-}$ is of type (1,k1,1,k1)(1,k-1,1,k-1) where $k=\operatorname{MIN}_{\operatorname{vert}}(S^{-})+1$
  5. (F5F{5})
    $\chi_{h}(\operatorname{Fan}(T))=\operatorname{\mathbf{i}}(S^{+},S^{-})$
  6. (F6F{6})
    $S^{+}$ is of type $(0,l+2,0,l)$ for some $l$.
  7. (F7F{7})
    $\operatorname{MIN}_{\operatorname{vert}}(S^{+})=\operatorname{MIN}_{-}(\operatorname{Fan}(T)^{+})$
  8. (F8F{8})
    $e(S^{-},S^{+})$ satisfies property ()(*) (Definitions 6.22 and 6.1).
Proof.
By Lemma 7.6 assume without loss of generality that $T$ is the $\operatorname{Fan}(T)$ of Lemma 7.6, i.e. $\operatorname{MIN}_{\operatorname{ray}}(T^{-})$ is realized by $J_{0}=[o,v^{+}]$. We will describe the process depicted in Figures 10(b)-(f). For any $\delta>0$ denote
$\tau(\delta)=[b_{+}(\delta),v^{+}-(\delta,0)]\times\left[\tfrac{1}{2}-\delta,\tfrac{1}{2}+\delta\right].$
By general position and centeredness of $T$, we can assume without loss of generality that for a sufficiently small $\delta$ the set $T\cap\tau(\delta)$ is a collection of vertical straight lines and the number of those lines equals $\operatorname{MIN}_{\operatorname{ray}}(T)$. Assume that $\delta$ is also so small that each of the (four) sets $\square_{\delta}\cap T^{\pm}$ and $U(\delta)\cap T^{\pm}$ is one line segment where $U(\delta)$ is as defined before the Straightening Lemma 7.5, see Figures 11(c) and 12(a). We will now construct a homeomorphism $\psi\colon\square\to\square$ which will realize the fan-move depicted on Figures 10(b)-(f). Let Let $\eta^{\prime}\colon\partial\square\to S^{1}$ be the homeomorphism given by $\eta^{\prime}(x)=(x-o)/|x-o|$ and let $\eta\colon\square\to D^{2}=\{(x,y)\in\mathbb{R}^{2}\,:\,x^{2}+y^{2}\leqslant 1\}$ be be the radial extension of $\eta^{\prime}$ such that $\eta(o)=\bar{0}$. Note that for all $\leqslant 1, r\partial\square_{r}$ maps to $\partial B(\bar{0},r)$.
[…]
Figure 11: Illustrating the maps $\psi$ and $\eta^{-1}\circ\psi\circ\eta$ from the proof of Lemma $7.$ Figure (a) is the domain and (b) the range of $\psi$. The colored regions are mapped respectively to the same colored regions. The light-red area is where $\psi$ is the identity. The yellow and orange areas are squeezed and the blue and green areas are stretched. The gray areas is where $\psi$ is defined by interpolating between the light-red and other colored areas. In (c) and (d) the domain and range of $\eta^{-1}\circ\psi\circ\eta$ is shown. The green and blue areas form an area inside of $\tau(\delta)$ according to (7.1). Figures (c) also shows the line segments $K^{+}\cap U(\delta)$ and $K^{+}\cap\square_{\delta}$ (red) as well as $K^{-}\cap U(\delta)$ and $K^{-}\cap\square_{\delta}$ (blue). In (d) the images of these line segments are shown. Notice how the gray area “drags” them into the lower region due to the interpolation.
For convenience, denote $\langle t,\theta\rangle=te^{i\theta}\in D^{2}$ (for $t\leqslant 1$). There is $\varepsilon>0$ such that
$\eta[\tau(\delta)]\supset\{te^{i\theta}\,:\,t\in[\delta,1-\delta],\theta\in[-\varepsilon,\varepsilon]\}.$ (7.1)
Let $\psi\colon D^{2}\to D^{2}$ be given by
[…]
Note that $\psi(\langle t,\theta\rangle)$ is not yet defined for $t\in[\delta/2,\delta]\cup[1-\delta,1-\delta/2]$. For these values of $t$, $\psi(t)$ will be defined by a “circular” interpolation as follows. For $t\in[\delta/2,\delta]$ let
$\psi(t,\theta)=\Big\langle t,\ (2-2t/\delta)\Theta\big(\psi(\langle\delta/2,\theta\rangle)\big)+(2t/\delta-1)\Theta\big(\psi(\langle\delta,\theta\rangle)\big)\Big\rangle$
and for $t\in[1-\delta,1-\delta/2]$ let
$\psi(t,\theta)=\Big\langle t,\ (2(1-t)/\delta-1)\Theta\big(\psi(\langle 1-1/8,\theta\rangle)\big)+(2(t-1)/\delta+2)\Theta\big(\psi(\langle 1-\delta/2,\theta\rangle)\big)\Big\rangle$
where $\Theta(\langle t,\theta\rangle)=\theta$. See Figure 11 for a visual explanation.
Consider the map $\eta^{-1}\circ\psi\circ\eta$. It performs the corresponding shift in $\square$ and stretches what used to be short vertical lines in $\tau$ in a fan-like manner to go around $o$ as shown in Figures 10(b)-(f), although we are still missing the final step. We should package the lower area into the rectangle $G=[1/4,3/4]\times[0,1/2-\delta]$. We will not explicitly construct this last homeomorphism $\zeta$ and will be content with Figure 12. Let also $g\colon G\to\square$ be the (unique) axis-aligned linear homeomorphism.
Both homeomorphisms $\eta^{-1}\psi\eta$ and $\zeta$ keep $\partial\square$ and a neighborhood of $o$ fixed, so $h=\zeta\eta^{-1}\psi\eta$ realizes an $R^{*}_{0}$-move. Let $\operatorname{Fan}(T)=\zeta\eta^{-1}\psi\eta T$. The final result without extra colors is depicted on Figure 13(e)(a). If $x$ is a crossing of $\operatorname{Fan}(T)$, then there is a crossing $y$ of $T$ such that $h(y)=x$. Since there are no crossings in $\tau\cup\square_{\delta}\cup U(\delta)$, $y$ must be such that if $\eta(y)=\langle t,\theta\rangle$, then $t\in[\delta,1-\delta]$ and $\theta\notin[-\varepsilon,\varepsilon]$ (yellow or orange areas in Figures 11(a),(c)). But then it is easy to check that $\psi(\eta(y))$ ends up in the area $\{\langle t,\theta\rangle\,:\,t\in[\delta,1-\delta],\theta\in[-3\pi/4,\pi/4]\}$ (yellow and orange in Fig. 11(b),(d)). This is in turn which is mapped by $\eta^{-1}$ to an area which is then squeezed into $G$ by $\zeta$. Thus, $x\in G$ which proves the first part of (F1F{1}). But $\operatorname{Cross}_{h}(T)\subset\operatorname{Cross}(T)$, so it proves also the second part. The fact that $\operatorname{MIN}_{\operatorname{ray}}(T)=\operatorname{MIN}_{\operatorname{ray}}(\operatorname{Fan}(T))$ follows from Lemma 5.2. By Lemma 7.3, $\operatorname{MIN}_{+}(\operatorname{Fan}(T)^{-})\geqslant\operatorname{MIN}_{\operatorname{ray}}(\operatorname{Fan}(T))$ so we need to prove $\operatorname{MIN}_{+}(\operatorname{Fan}(T)^{-})\leqslant\operatorname{MIN}_{\operatorname{ray}}(T)$. Let $J$ be the vertical line $o,o^{+}]$. It is clear from the construction that $h^{-1}J$ intersects $T^{-}$ exactly in the vertical lines of $T^{-}\cap\tau$. In fact $h^{-1}J\subset\square_{\delta}\cup J_{0}\cup U(\delta)$ and $U(\delta)\cap T^{-}\cap h^{-1}J=\varnothing$ and $\square_{\delta}\cap T^{-}\cap h^{-1}J=\{o\}$. Here $J_{0}$ is from the very beginning of this proof. Thus, $\operatorname{\mathbf{i}}(J,\operatorname{Fan}(T)^{-})=\operatorname{\mathbf{i}}(h^{-1}J,T^{-})=\operatorname{\mathbf{i}}(J_{0},T^{-})=\operatorname{MIN}_{\operatorname{ray}}(T^{-})$. This proves (F2F{2}). Before proving (F3F{3}), let us note that the areas in the vertical edges of $G$ which intersect $T^{-}$ are images under $h$ of the horizontal edges of $\tau$. On the other hand the horizontal edges of $G$ intersect $T^{-}$ each in one point. The lower edge at $o^{-}$ and the top edge because it is the image of the lower edge of $\square_{\delta}$ under $h$. It therefore follows that
$S^{-}\text{ is of type }(1,\operatorname{MIN}_{\operatorname{ray}}(T^{-}),1,\operatorname{MIN}_{\operatorname{ray}}(T^{-}))$ (7.2)
[…]
$\operatorname{Fan}(T)^{+}\cap\operatorname{Fan}(T)^{-}=\Big(\operatorname{Fan}^{\circ}(T)^{+}\cap\operatorname{Fan}^{\circ}(T)^{-}\Big)\cup\{o\},$
but $o\notin G$, so
$\operatorname{Fan}(T)^{+}\cap\operatorname{Fan}(T)^{-}\cap G=\operatorname{Fan}^{\circ}(T)^{+}\cap\operatorname{Fan}^{\circ}(T)^{-}\cap G$
and therefore, by (F1F{1}), $\operatorname{Cross}_{h}(\operatorname{Fan}(T))=\operatorname{Fan}(T)^{+}\cap\operatorname{Fan}(T)^{-}\cap G$ We now have:
$\operatorname{\mathbf{i}}(S^{+},S^{-})$ $=\operatorname{Card}(g(\operatorname{Fan}(T)^{+}\cap G)\cap g(\operatorname{Fan}(T)^{-}\cap G))$ $\text{by definition of }S^{\pm}\text{ and the above}$
$=\operatorname{Card}((\operatorname{Fan}(T)^{+}\cap G)\cap(\operatorname{Fan}(T)^{-}\cap G))$ $\text{because }g\text{ is a homeomorphism}$
$=\operatorname{Card}(\operatorname{Fan}(T)^{+}\cap\operatorname{Fan}(T)^{-}\cap G)$
$=\operatorname{Card}(\operatorname{Cross}_{h}(\operatorname{Fan}(T)))$ by the above
$=\chi_{h}(\operatorname{Fan}(T))$
For (F6F{6}), let $l$ be the number of lines in $T^{+}\cap\tau$. From the construction it follows that this is the same as the number of lines of $\operatorname{Fan}(T)^{+}$ entering $G$ through the right vertical edge. On the left vertical edge two additional lines of $\operatorname{Fan}(T)^{+}$ come into $G$ from the areas $h\square_{\delta}$ and $hU(\delta)$ (see Figure 12(c),(d) and the caption of that figure). No $\operatorname{Fan}(T)^{+}$ lines intersect $G$ at the horizontal edges. This proves (F6F{6}).
Let $J_{v}$ be a vertical splitting curve for $S^{+}$ with $i(J_{v},S^{+})=\operatorname{MIN}_{\operatorname{vert}}(S^{+})$. Then $g^{-1}J$ can be modified in the same manner as in the proof of (F3F{3}) above to be connected to $o$ without introducing new intersections with $\operatorname{Fan}(T)^{+}$. Further, it can be modified within $U(\delta)\cap\square^{-}$ to be connected to $o^{-}$ still without new intersections with $\operatorname{Fan}(T)^{+}$ (recall that $(U(\delta)\setminus T^{+})\cup\{o^{-}\}$ is path-connected by the choice of $\delta$, and so is this set intersected with $\square^{-}$). This shows that
$\operatorname{MIN}_{\operatorname{vert}}(S^{+})\geqslant\operatorname{MIN}_{-}(\operatorname{Fan}(T)^{+}).$ (7.3)
Suppose that $J^{-}$ is a lower splitting curve for $T^{+}$ such that $\operatorname{\mathbf{i}}(J^{-},\operatorname{Fan}(T)^{+})=\operatorname{MIN}_{-}(\operatorname{Fan}(T)^{+})$. If $g(J^{-}\cap G)$ is a vertical splitting curve, then
$\operatorname{MIN}_{\operatorname{vert}}(S^{+})$ $\leqslant\operatorname{\mathbf{i}}(g(J^{-}\cap G),S^{+})$
$=\operatorname{\mathbf{i}}(J^{-}\cap G,T^{+}\cap G)$ $\text{because }g\text{ is homeo and def. of }S^{+}$
$\leqslant\operatorname{\mathbf{i}}(J^{-},T^{+})$ $\text{because }J^{-}\cap G\cap T^{+}\cap G\subset J^{-}\cap T^{+}.$
$=\operatorname{MIN}_{-}(\operatorname{Fan}(T)^{+})$ $\text{by the choice of }J^{-}.$
Otherwise, note that there is a homeomorphic transformation $\varphi$ of $\square^{-}$ into its subset such that any lower splitting curve is mapped onto a curve whose intersection with $G$ is a vertical splitting curve of $G$, and such that $\varphi[\operatorname{Fan}(T)^{+}]=\operatorname{Fan}(T)^{+}\cap G$. The construction of $\varphi$ is illustrated in Figure 13(e). Let $J=g^{-1}\varphi J^{-}$. Then by the construction of $g$ (Figure 13(e)) $J$ is a vertical splitting curve for $S^{+}$ and we have
$(g^{-1}\varphi J^{-})\cap S^{+}=\varphi J^{-}\cap\operatorname{Fan}(T)^{+}\cap G=J^{-}\cap\varphi^{-1}\operatorname{Fan}(T)^{+}=J^{-}\cap\operatorname{Fan}(T)^{+},$
so $\operatorname{\mathbf{i}}(J,S^{+})=\operatorname{\mathbf{i}}(J^{-},\operatorname{Fan}(T)^{+})=\operatorname{MIN}_{-}(\operatorname{Fan}(T)^{+})$, so it again follows that $\operatorname{MIN}_{\operatorname{vert}}(S^{+})\leqslant\operatorname{MIN}_{-}(\operatorname{Fan}(T)^{+})$. Together with (7.3) this proves (F7F{7}).
It remains to prove (F8F{8}), so let us show that $e(S^{-},S^{+})$ satisfies property ()(*) (Definition 6.1). To be consistent with Definition 6.22, denote $T=S^{-}$ and $S=S^{+}$. Let $S^{*}$ be a smoothing of $S$ as in Definition 6.22. Then endpoints of $S^{*}$ are the same as the endpoints of $S$. Recall the notation $L_{i}(T)$ (Definition 4.6). These are open subarcs of the left and right edge of $\square$ separated by the endpoints of $T=S^{-}$. Via $g^{-1}$ they become subarcs of vertical edges of $G$. Since there are k1k-1 endpoints of $T^{-}$ no each side of $G$ by ((F4F{4})), there are $k$ such subarcs on each side. Let $c_{i}$ be the number of endpoints of $S=S^{+}$ in $L_{i}(T)$. Recall that the sets $L_{1}(T),\dots,L_{n}(T)$ ($n=2k$) are enumerated counterclockwise starting from the top left. Since the arcs that go around in the fan correspond to the original lines in $\tau\cap T$, each point $S\cap L_{i}(T)$ is matched by a point in $S\cap L_{f(i)}(T)$ where f(i)=ni+1f(i)=n-i+1 as defined in the beginning of Section 6.1. Figures 12(d), 13(e)(a) as well as 10(f) illustrate this matching via red arcs. Therefore, if $i\in\llbracket 2,k-1\rrbracket$, we have $c_{i}=c_{f(i)}$. But $c_{i}=s_{i}(e(T,S))$, so this proves ($enumi_{1}$). When $k>1$, there is one extra endpoint of $S$ in $L_{1}(T)$ corresponding to the arc of $T^{+}$ which originates in $o$ and one extra endpoint in $L_{k}(T)$ corresponding to the arc which originates in $o^{+}$ (red arcs in the figures). This proves ($enumi_{2}$). Clearly, if an arc goes from $L_{i}(T)$ to $L_{j}(T)$, the same arc goes from $L_{j}(T)$ to $L_{i}(T)$, so the matrix is symmetric verifying $e\in\mathcal{M}^{s}_{n}(\mathbb{N})$ and condition ($enumi_{3}$) follows immediately from Definition 6.22 of $e(T,S)$. ∎
Figure 12: (a): Same as Figure 11(c) but with the vertical lines of $K\cap\tau$ shown. (b): Same as Figure 11(c) but with the stretched lines shown. Notice how the gray area “drags” the two red lines into the lower region due to the nature of the interpolation. (c): The result of the application of the homeomorphism $\zeta$ which “packages” the lower part of the knot into the rectangle $G$. (d): The rectangle $G$ is overlayed in white. All the properties (F1F{1})(F8F{8}) of the Fan Lemma 7.7 can be “read” from Figure (d). For example the number of blue lines (lines of $K^{-}$) which go around the “fan” equals to $\operatorname{MIN}_{\operatorname{ray}}(K^{-})=\operatorname{MIN}_{\operatorname{ray}}(\operatorname{Fan}(K)^{-})$ because it is the same as the number of blue lines shown in (a) to the right of the small gray square (denoted $\square_{\delta}$ in the proof). The “+2” in (F6F{6}) comes from the two red lines which arrive into $G$ from the gray areas.
Figure 13: (a): The result of the procedure depicted in Figures 11 and 12. Figures (b)-(e) show the construction of the homeomorphism $\varphi$ from the proof of Lemma 7.7(F7F{7}). (b): The light-gray area is $\square^{-}$ which needs to be transformed such that any lower splitting curve becomes a vertical splitting curve in $G$. We have added an example lower splitting curve $J$ in green color. It goes partially in $G$ and partially outside of it, but necessarily in $\square^{-}$ by definition. Note also that by definition it must start in $o$ and end in $o^{-}$. (c): First move the red curve (part of $K^{+}$) coming from $o$ a bit to clear the top area. This of course forces $J$ to move too. We have also chosen $\varepsilon>$ such that in the $\varepsilon$-neighborhood of the vertical edges of $G$, $\operatorname{Fan}(K)$ is still trivial and consists of horizontal continuations of the lines that come into $G$ from the sides. At this point we mark with darker green those parts of $J$ which will never need to move for $\varphi$ will be identity in “deep” inside $G$. (d): Start the transformation by modifying $\square^{-}$ from the sides while keeping the curves intact (i.e. $\varphi[K]\subset K$), but squeezing $J$. (e): In the end $J$ is squeezed entirely into $G$ and $\square^{-}$ is squeezed into a rectangle whose vertical sides contain the vertical sides of $G$. Just like our green $J$, every lower splitting curve must split $G$ vertically. The constructed homeomorphism $\varphi$ keeps the rectangle $1/4+\varepsilon,3/4-\varepsilon]\times[0,1/2]$ fixed with possible exceptions above $G$ (in the beginning of the construction). In this case the number of intersection between $J$ and $K^{+}$ (red) outside $G$ before the transformation is 4. All these intersections are moved inside $G$ by the end of the transformation.
The following is the main result of this section.
Proposition 7.1 (A Bound for Hybrid Crossings for Tangle Diagrams).
Suppose $T$ is a centered subdivided t-diagram. Then
$\operatorname{\mathbf{i}}(T^{-},T^{+})\geqslant\operatorname{MIN}_{\operatorname{ray}}(T^{-})+\operatorname{MIN}_{\operatorname{ray}}(T^{+}).$
Proof.
Let $\operatorname{Fan}(T)$ be the subdivided tangle diagram given by Lemma $7.$ Since $\operatorname{Fan}(T)\sim^{*}_{0}T$ and $\operatorname{MIN}_{\operatorname{ray}}$ is invariant under $\sim^{*}_{0}$ (Lemma 5.2 applies to composite tangle as well as knot diagrams), we can assume without loss of generality that $\operatorname{Fan}(T)=T$. Let $S^{-}$ and $S^{+}$ be as defined in the statement of Lemma $7.$ By (F4F{4}), (F6F{6}), and (F8F{8}), $S^{-}$ and $S^{+}$ satisfy the assumptions of Proposition 6.1, so it follows that
$\chi_{h}(\operatorname{Fan}(T))$ $=\operatorname{\mathbf{i}}(S^{+},S^{-})$ by (F5F{5})
$=\operatorname{Card}(S^{-}\cap S^{+})$ $\text{by the definition of }\operatorname{\mathbf{i}}(\cdot,\cdot)$
$\geqslant\operatorname{MIN}_{\operatorname{vert}}(S^{+})+k-1.$ by Proposition 6.1
$=\operatorname{MIN}_{\operatorname{vert}}(S^{+})+\operatorname{MIN}_{\operatorname{vert}}(S^{-})$ by the definition of $k$ in (F5F{5})
$=\operatorname{MIN}_{+}(\operatorname{Fan}(T)^{-})+\operatorname{MIN}_{-}(\operatorname{Fan}(T)^{+})$ by (F3F{3}) and (F7F{7})
[…]
Corollary 7 (A Bound for Hybrid Crossings of a Knot Diagram).
Suppose $K$ is a centered knot diagram. Then $\operatorname{\mathbf{i}}(K^{-},K^{+})\geqslant\operatorname{MIN}_{\operatorname{ray}}(K^{-})+\operatorname{MIN}_{\operatorname{ray}}(K^{+})$.
Proof.
Let $T=(K,K^{+},K^{-},\{1\},\varnothing,\ell)$. Then $T$ is a centered subdivided t-diagram. The result follows from Proposition 7.1. ∎
8 Proof of Theorem 3.1
[…]
Proof.
[…]
$\geqslant\chi(\widehat{K}^{+})+\chi(\widehat{K}^{-})+\operatorname{MIN}_{\operatorname{ray}}(\widehat{K}^{-})+\operatorname{MIN}_{\operatorname{ray}}(\widehat{K}^{+})$ by Corollary 7
$=\chi(\widehat{K}^{+})+\operatorname{MIN}_{\operatorname{ray}}(\widehat{K}^{+})+\chi(\widehat{K}^{-})+\operatorname{MIN}_{\operatorname{ray}}(\widehat{K}^{-})$ rearranging terms
[…]
9 Proof of Theorem 3.3
[…]
Proof.
By using subdivided tangle diagrams instead of knots (Definition 7.1) one straightforwardly adapts all the proofs of Section 5 to links, including Corollary 5. Then the proof is the same as that of Theorem 3.1 presented in Section 8. (Use Proposition 7.1 instead of Corollary 7.) ∎
10 Discussion and Open Questions
Question 1.
Is there a less ad hoc topological interpretation (than that presented in Corollary 2) of the fact that it is impossible to reduce the number of crossings of a composite knot using moves in $\mathcal{R}^{*}_{opp1}$?
Question 2.
Find invariants of knot diagrams $^{-}(K)$ and $^{+}(K)$ such that Corollary 5 can be proved with $^{\pm}(K)$ instead of $\operatorname{MIN}_{\operatorname{ray}}(K^{\pm})$ and with $\sim_{\mathcal{R}}$ instead of $\sim^{X}_{\mathcal{R}}$ but also such that Proposition 7.1 still holds when $^{\pm}(K)$ is used instead of $\operatorname{MIN}_{\operatorname{ray}}(K^{\pm})$. This would resolve the addivitiy of the crossing number conjecture by an analogue of the proof of Theorem 3.1.